### v2

# Mutatis Mutandis: Revisiting the Comparator in Discrimination Testing

Abstract.
Testing for discrimination consists of deriving a profile, known as the comparator, similar to the profile making the discrimination claim, known as the complainant, and comparing the outcomes of these two profiles. An important aspect for establishing discrimination is evidence, often obtained via discrimination testing tools that implement the complainant-comparator pair. In this work, we revisit the role of the comparator in discrimination testing. We argue for the causal modeling nature of deriving the comparator, and introduce a two-kinds classification for the comparator: the ceteris paribus (CP), and mutatis mutandis (MM) comparators. The CP comparator is the standard one among discrimination testing, representing an idealized comparison as it aims for having a complainant-comparator pair that only differs on membership to the protected attribute. As an alternative to it, we define the MM comparator, which requires that the comparator represents what would have been of the complainant without the effects of the protected attribute on the non-protected attributes. The complainant-comparator pair, in that case, may also be dissimilar in terms of all attributes. We illustrate these two comparators and their impact on discrimination testing using a real illustrative example. Importantly, we position generative models and, overall, machine learning methods as useful tools for constructing the MM comparator and, in turn, achieving more complex and realistic comparisons when testing for discrimination.
algorithmic fairness, discrimination discovery, counterfactual model of discrimination
1. Introduction
The concept of discrimination refers to an unjustified difference in treatment toward an individual or group of individuals based on (perceived) membership to a protected-by-law group (Romei and Ruggieri, 2014). We view, thus, discrimination as unfairness sanctioned by law. Testing for discrimination against an individual instance is a comparative process centered around the complainant-comparator pair. As the names would suggest, the complainant is the individual instance motivating the discrimination claim in court, while the comparator is the individual instance used for validating the claim. We use the comparator to answer the hypothetical question behind any discrimination claim: would the decision have been the same if the complainant had been a member of the non-protected group instead of the protected group under non-discrimination law? The comparison between the outcomes of the complainant and its comparator helps establish whether the decision-making process in question is or not discriminatory. Most tools for discrimination testing, from the traditional methods like correspondence studies (Bertrand and Mullainathan, 2004) and situation testing (Bendick, 2007) to the recent algorithmic methods like discrimination discovery (Ruggieri et al., 2010) and counterfactual fairness (Kusner et al., 2017), rely on this pair. Importantly, these tools not only implement the complainant-comparator comparison but also derive the complainant’s comparator as a whole.
Non-discrimination law requires similar, or similarly situated, individuals to be treated similarly regardless of gender, race, or other protected attributes. It is a requirement largely shared across countries, only with some variation on the list of protected attributes (Romei and Ruggieri, 2014). The comparator is meant to represent a similar (or, at least, comparable) individual to the complainant, allowing to envision an alternative path for the complainant conditional on “changing” its protected attribute. It requires what we view, in practice, as an imaginative act in which we find (e.g., via matching models (Qureshi et al., 2020)) or create (e.g., via generative models (Black et al., 2020)) an individual instance representative enough of the complainant’s alternative world. In fact, the comparator, unlike the complainant, can also be a fictitious individual instance (Weerts et al., 2023). In a time of modern decision flows in which machine learning (ML) methods are increasingly used to aid as well as regulate decision-making processes, such modeling flexibility around the comparator offers a promising ML application.
The choice of comparator, meaning how these tools for discrimination testing derive the comparator, matters for three incremental reasons. First, it embodies a normative statement on what similarity between individuals is and, in turn, how the protected attribute influences individual outcomes within the context of interest (e.g., (Weerts et al., 2023)). Second, it conditions the results of these tools by determining which individual outcomes are to be compared for evaluating the discrimination claim (e.g., (Álvarez and Ruggieri, 2023)). Third, it acts as a target for the sort of substantive equality between individuals that we wish to achieve as a society through non-discrimination law (e.g., (Wachter et al., 2020)). Most tools aim to define the comparator as an observably similar instance to the complainant all but for membership to the protected attribute, such as defining a male candidate with the same work experience as the female complainant in a hiring decision or a white applicant with the same credit score as the non-white complainant in a lending decision. We refer to this standard operationalization of the comparator as the ceteris paribus, or “with all else being equal,” (CP) comparator. Precisely because of these reasons, debates around the CP comparator predate modern decision flows, e.g., (Westen, 1982; Heckman, 1998; Kohler-Hausmann, 2019), and remain relevant today as fair ML researchers tackle the new problem of algorithmic discrimination, e.g., (Kleinberg et al., 2019; Ruggieri et al., 2023; Álvarez et al., 2024). We would add that this reliance on the CP comparator is problematic for fair ML. It overlooks the potential positive role ML methods, particularly generative models, could play as tools for deriving more complex comparators that address the ongoing debates.
This work studies the problem of testing for discrimination by revisiting the role of the comparator. We do so by, first, embracing the causal modeling nature of discrimination testing and, second, proposing an alternative to the CP comparator. Regarding the first point, given that the comparator answers to the “what would have been if” question around the complainant, we assert that the comparator is a counterfactual representation of the factual complainant. We further assert that the derivation of the comparator itself constitutes a causal modeling problem and that such problem formulation is, for now, the best possible one given the comparative nature of discrimination testing. Regarding the second point, we propose the mutatis mutandis, or “with the appropriate adjustments being made,” (MM) comparator. The MM comparator requires that all non-protected attributes constituting the comparator account for the downstream effects of the protected attribute. The CP comparator, thus, represents an idealized comparison to the complainant, while the MM comparator represents an adjusted comparison to the complainant. With the MM comparator, we expand the problem of discrimination testing and, given the greater complexity of deriving a MM comparator relative to a CP comparator (read, a stronger “imagination”), we identify an impactful application for fair ML models. Example 1.1 below motivates and illustrates the difference between the two comparators.
Example 1.1.
(Tenure at U) This example is based on Morgan et al. (2021)’s research on the unequal impact of parenthood in academia for female researchers. [^1] Social expectations and household dynamics are key factors in creating this difference. Mothers, as the primary caregivers, tend to carry the burden of childbearing within their households, which consumes potential time that could be spent on advancing ones academic career by, e.g., writing papers at home or networking at conferences. Fathers tend to share less of this burden. See Morgan et al. (2021) for details. Also see Roman and Rodríguez (2023). Let us assume that the university U only considers the number of publications for granting tenure. Clara, a female researcher, was denied tenure for having published only 12 papers and files a discrimination claim against U, becoming the complainant. Suppose that we use Mike, a male researcher that also published 12 papers and was denied tenure by U, as her comparator. Do we deny Clara’s discrimination claim? Given what we know about parenthood’s impact on academic performance, would our answer change by considering that Clara is a mother? Assuming some quantifiable penalty on the number of publications from motherhood, would we be willing to use Vincent, with 18 publications and recently tenured, as Clara’s comparator instead of Mike?
In Example 1.1 Mike represents the CP comparator (Definition 2.1) while Vincent the MM comparator (Definition 3.1). Depending on which one we use, assuming a literal comparison, we would invalidate or validate, respectively, Clara’s discrimination claim. Despite its simplicity, Example 1.1 captures the intricacies behind discrimination testing once we contextualize the decision-making process based on information about the protected attribute. Given what we know from works like Morgan et al. (2021), we are implicitly asking which male researcher is more similar to Clara? It is not an easy question to answer, highlighting the complexity behind modeling the comparator. With Example 1.1 and, overall, with this work we are not suggesting that the CP comparator is wrong or that the MM comparator is better. Instead, we want to show that the MM comparator is an alternative formulation to the CP comparator and, overall, stress the importance of the choice of comparator. Policies like non-discrimination law and positive actions aimed at addressing/removing systematic inequalities require concrete targets (Wachter et al., 2020). Discrimination testing tools, essentially, enforce such targets through the comparator, with the CP and MM comparators representing two distinct targets.
Structure.
In the remainder of this section we discuss our modeling choice for causality (Section 1.1) and relevant related work (Section 1.2). Moving forward, in Section 2 we argue that deriving the comparator is a causal modeling problem and formulate the CP comparator, which is the standard in discrimination testing. In Section 3, we propose the MM comparator and justify its departure from the CP comparator. In Section 4, using a popular real dataset in fair ML, we showcase both comparators by comparing two related discrimination testing tools—situation testing (Thanh et al., 2011) and counterfactual situation testing (Álvarez and Ruggieri, 2023)—that implement, respectively, the CP comparator and MM comparator. In Section 5, we provide a representative survey of discrimination testing tools. In Section 6, we conclude.
Contributions.
Our main contributions are threefold. First, we demonstrate why the derivation of the comparator is best understood as a causal modeling problem. Second, with the MM comparator we propose a new approach to deriving the comparator and, in turn, introduce the distinction between CP and MM comparators. Our experiments illustrate how this distinction impacts discrimination testing. Third, we position ML methods as a promising tool for achieving the MM comparator and developing better tools for discrimination testing.
Audience and scope.
Discrimination testing is an interdisciplinary topic. Our work is aimed primarily at ML researchers working on this topic and, hence, our framing of it as a modeling problem. We are interested in quantitative tools used for deriving prima facie evidence for discrimination claims, which constitutes one aspect of the legal process for uncovering discrimination. Our focus is mainly on the operationalization of the comparator, which we view as a ML problem that holds for any modern decision flow, regardless of the nature of the decision-maker. Throughout the paper, we use generative model to refer to a model capable of creating synthetic data (McElreath, 2018). Similarly, we use counterfactuals to refer to interventional probability distributions obtained through causal reasoning (Pearl, 2009), not to be confused with counterfactual explanations222 Counterfactual explanations refer to individual instances similar to the one that is explained but have a different outcome. To avoid confusion, Miller (2019) proposes to name them contrastive explanations, and reserve the term counterfactual as understood in the statistical causality literature. from explainable AI (Stepin et al., 2021; Guidotti, 2024).
1.1. Auxiliary Causal Knowledge
We formulate causality using structural causal models (SCM) (Pearl, 2009), which is the most used framework within ML. SCM are probabilistic graphical models that link (latent) exogenous variables to (observable) endogenous variables via structural equations. Each SCM describes a probability distribution and can generate hypothetical distributions via interventions. Formally, a SCM $\mathcal{M}$ is a tuple $\mathcal{M}=\langle\mathbf{U},\mathbf{X},\mathbf{F}\rangle$, where $\mathbf{U}=(U_{1},\ldots,U_{p})$ are $p$ independent exogenous latent variables, with $\mathbf{U}\sim P(\mathbf{U})=\prod_{j}P(U_{j})$, $\mathbf{X}=(X_{1},\ldots,X_{p})$ are $p$ endogenous observed variables, and $\mathbf{F}=(f_{1},\ldots,f_{p})$ are structural equations such that:
(1) $\text{for $j=1,\dots,p$}\;\;\;X_{j}:=f_{j}(X_{pa(j)},U_{j})$
Each $f_{j}$ defines the observed variable $X_{j}$ on the basis of its causal parents $X_{pa(j)}$ and of the latent variable $U_{j}$ Associated to the SCM $\mathcal{M}$ is the causal graph $\mathcal{G}$, in which each node is a random variable and the directed edges between them a causal effect. We assume a directed acyclical graph, or DAG, $\mathcal{G}$. Through $\mathcal{G}$ we represent how the variables relate to each other in terms of cause-effect pairs. For instance, the causal graph $X_{1}\rightarrow X_{2}$ reads as “$X_{1}$ causes $X_{2}$”.
The SCM $\mathcal{M}$ in (1) can act as a generative model. In particular, it can generate counterfactual distributions that answer to counterfactual questions as the ones asked in discrimination claims. Let $A$ be one of the variables in $\mathbf{X}$ denoting the protected attribute. We calculate the counterfactual variables $\mathbf{X}^{CF}_{A\leftarrow\alpha}$ resulting from setting $A:=\alpha$ through a three-step procedure consisting of abduction, action, and prediction (Pearl et al., 2016). Abduction: for each prior distribution $P(U_{j})$ that describes $U_{j}$, we compute its posterior distribution given the evidence, or $P(U_{j}\;|\;\mathbf{X})$. Action: we intervene on $A$ by changing its structural equation to $A:=\alpha$, which gives way to a new SCM $\mathcal{M}^{\prime}$. Prediction: we generate the counterfactual distribution $P(\mathbf{X}^{CF}_{A\leftarrow\alpha}\;|\;\mathbf{X})$ by propagating the abducted $P(U_{j}\;|\;\mathbf{X})$ through the revised structural equations in $\mathcal{M}^{\prime}$.
Beyond Section 4, in which we rely on counterfactual generation to implement the MM comparator, what we discuss in Sections 2 and 3 is not exclusive to SCM. It is our generative model of choice, mainly because previous works on discrimination testing (e.g., (Kusner et al., 2017; Kilbertus et al., 2017; Álvarez and Ruggieri, 2023)) use it. Here, our goal is not prioritize one generative model over the others, but to focus on one generative model to show how the others could be useful for the problem of discrimination testing.
1.2. Relevant Related Work
The literature on discrimination testing is vast, multidisciplinary, and regional. We refer to (Chen et al., 2024) for a recent survey on testing fairness of ML models, and to (Hutchinson and Mitchell, 2019; Romei and Ruggieri, 2014) for surveys with a multidisciplinary perspective. For a general overview of bias and fairness in AI, see (Álvarez et al., 2024; Ntoutsi et al., 2020; Mehrabi et al., 2022). Throughout the paper, we consider individual discrimination. Namely, we are interested at testing whether an individual instance of a protected group has been discriminated. The notion of group discrimination tackles instead discrimination against a subset of individuals from a protected group. However, Binns (2020) argues that the apparent conflict between individual and group fairness is an artifact of fairness metrics rather than reflecting different normative principles.
In this subsection, we position our work relative to three recent papers that we view as directly applicable to the modeling problem of revisiting the comparator. To the best of our knowledge, we are the first to explicitly tackle this problem. Later in Section 5, we provide a representative survey of discrimination testing tools, with a focus on the comparator chosen by those tools. We do so as we first want to introduce the CP and MM comparators (Sections 2 and 3), and illustrate their impact on a discrimination testing problem (Section 4).
First, Kohler-Hausmann (2019) argues that discrimination testing focuses on answering counterfactual questions about the complainant. She refers to the current approach as the causal counterfactual model of discrimination, and states that it is flawed due to its treatment of protected attributes. In particular, using a constructivist approach to race, she argues that comparing similar individuals that only differ on skin color fails to capture how race affects all other aspects of these individuals’ lives. We draw considerably from Kohler-Hausmann (2019) as she writes the first critique to the CP comparator. However, different from Kohler-Hausmann (2019), we propose an alternative to the CP comparator in the form of the MM comparator and embrace the role of counterfactual reasoning within discrimination testing. This last point is important as Kohler-Hausmann’s view (see, e.g., Hu and Kohler-Hausmann (2020)) is largely against causal modeling for discrimination testing and is suspicious toward attempts to measure the so-called causal effect of race. She fails, though, to provide an alternative to the current model. We disagree with her position toward causal modeling. What we agree with her is that we can (and, potentially, should) do better than the CP comparator. The MM comparator is a response to Kohler-Hausmann (2019).
Second, Wachter et al. (2020) state that non-discrimination law aims for substantive equality, meaning it should aim at changing the status quo of the societal context of interest. [^3] Their argument is specific to the European Union. Although it seems not to be the case for US non-discrimination law (Barocas and Selbst, 2016), the goal of substantive equality is implicit in many of the fair ML definitions used by US researchers as argued by Wachter et al. (2020). Under this premise, they classify fair ML methods into bias preserving and bias transforming based on whether they assume (the former) or not (the latter) a neutral status quo. Although Wachter et al. (2020) make a strong case for ML methods as societal shapers, they lack clarity on how these methods are to be implemented, specially, with respect to non-discrimination law. We tackle this gap by formalizing the MM comparator. Our reading of Wachter et al. (2020) is that non-discrimination law must not aim at reaching equality of outcomes only, but also at imposing equality of opportunities in a systematic way. Such a step requires envisioning an ideal target, which needs to be generated, for the current context to aim at. The MM comparator represents the goal of substantive equality in the context of discrimination testing.
Third, Álvarez and Ruggieri (2023) introduce a similar classification to the CP and MM comparators distinction. Drawing from Kohler-Hausmann (2019)’s discussion on the causal counterfactual model of discrimination, they state that discrimination testing tools can be broadly classified into comparators that do and do not consider the downstream effects of the protected attribute on all other attributes. They offer a valuable first step to the goal of revisiting the comparator, but fall short in formalizing their claims on the comparator. They neither define these kinds of comparators nor they survey the quantitative tools to illustrate them. On top of addressing these shortcoming from Álvarez and Ruggieri (2023), we discuss the role of counterfactual reasoning in discrimination testing, which they take for granted.
We note that the causal treatment of discrimination, specifically, using SCM is a common practice. See (Makhlouf et al., 2024) for a survey. In particular, Kilbertus et al. (2017) are the first to formulate problems like direct, indirect, and proxy discrimination using SCM, while Plecko and Bareinboim (2024) try to summarize discrimination within a general causal fairness model using SCM. Works like these are relevant, but lack interdisciplinary depth. For instance, both papers fail to discuss non-discrimination law and how exactly the modeling tools based on their formulations can be (or need to be) used as evidence. Their treatment of the subject is detached from the real-world application. These papers, e.g., do not discuss the comparator. To be clear, such papers have a different scope—mainly to formulate discrimination using SCM—but, because of such approach, fail to address the role of causality in discrimination testing. [^4] Such a discussion is more explicit in the social sciences, in which causal models are also used for testing discrimination. See, e.g., Heckman (1998). In that sense, our work complements these and similar works by revisiting the comparator.
2. The Comparator as a Causal Modeling Problem
Testing for discrimination amounts to a two-step modeling problem centered on deriving the comparator $i^{\prime}$, and comparing it to the discrimination claim’s complainant $i$. In this section, we focus on the derivation of the comparator, motivating the need for $i^{\prime}$ and its inherent challenges. Importantly, we argue that the derivation of the comparator is a causal modeling problem not only because we use the comparator to demonstrate whether the decision is or not caused by the protected attribute, but also because $i^{\prime}$ is a counterfactual representation of the factual $i$.
2.1. The Setup
Non-discrimination law mandates that a decision-maker—be it algorithmic or human—treats similar individuals similarly regardless of the protected attribute (Kleinberg et al., 2019). Discrimination testing tools uncover discrimination by finding and comparing similar individuals to isolate the effect of the protected attribute on the decision outcome (Romei and Ruggieri, 2014). Formally, letting $A$ represent the protected attribute, $\mathbf{X}$ the set of non-protected attributes, and $Y$ the decision outcome, we write this two-step modeling problem into Definitions 2.1 and 2.2, respectively.
Definition 2.1.
(The CP Comparator) Let the individual profile $\langle\mathbf{x}_{i},a_{i},y_{i}\rangle$ represent the complainant $i$ on which the discrimination claim is based. We define a ceteris paribus (CP) comparator $i^{\prime}$ as an individual profile $\langle\mathbf{x}_{i^{\prime}},a_{i^{\prime}},y_{i^{\prime}}\rangle$ such that:
(2) $d\big{(}\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\big{)}\leq\epsilon\;\;\;\text{and}\;\;\;a_{i}\neq a_{i^{\prime}}$
where $d$ is a distance function [^5] The issue of engineering fairness is, without doubts, challenging (Scantamburlo, 2021). In this regard, we believe that the definition of the distance function $d$ is not only a technical choice (see the survey (Blanco-Mallo et al., 2023)), but is necessarily specific of the application domain at hand. For instance, in job screening, $d$ can result from a weighting of the skills of job candidates: how to set the weights is a non-trivial issue, which requires an agreement among multiple stakeholders. For this reason, in this paper, we make no assumption/preference about the definition of $d$. (the closer to 00 the more similar are $\mathbf{x}_{i}$ and $\mathbf{x}_{i^{\prime}}$) and ϵ+italic-$\epsilon\in\mathbb{R}^{+}$ a threshold.
Importantly, it is possible to provide a hypothetical comparator (Weerts et al., 2023), meaning we may find or generate an $i^{\prime}$ similar under $d$ to $i$. The comparator derived from (2) is the standard comparator in discrimination testing (Heckman, 1998; Kohler-Hausmann, 2019). [^6] There is a conception of discrimination testing that is non-comparative, though it is not widely used. See Binns (2020). We refer to it as the ceteris paribus, or “with all else being equal,” (CP) comparator as, ideally, we expect to find an $i^{\prime}$ profile all equal but for the protected attribute to the $i$ profile when $\epsilon=0$. Intuitively, it operationalizes the basic axiom underpinning non-discrimination law—that of treating similar individuals similarly—especially when the decision-maker $f$ is, realistically, only allowed to use the set of non-protected attributes when making a decision, meaning $Y=f(\mathbf{X})$. In Section 3 we come back to the CP comparator and revisit it explicitly.
Definition 2.2.
(Difference in Outcomes) Given the individual profiles $\langle\mathbf{x}_{i},a_{i},y_{i}\rangle$ and $\langle\mathbf{x}_{i^{\prime}},a_{i^{\prime}},y_{i^{\prime}}\rangle$, we test the discrimination claim by focusing on the difference in outcomes:
(3) $\delta=P(Y=y_{i^{\prime}}\,|\,\mathbf{X}=\mathbf{x}_{i^{\prime}},A=a_{i^{\prime}})-P(Y=y_{i}\,|\,\mathbf{X}=\mathbf{x}_{i},A=a_{i})$
where, $\delta$ may, e.g., be greater or equal to 0 or need to follow a known distribution depending on the legal context.
The difference in outcomes (3) is not based on a literal comparison, meaning $y_{i^{\prime}}-y_{i}$, but on a difference in probabilities. It implies the need to construct some measure of uncertainty (hence, why we write $\delta$ in terms of probabilities) around the decision outcome for each individual profile. Formally, $\delta$ is the difference between two point estimates that requires the use of a statistical estimator. To obtain it, therefore, we need either multiple observations centered on these two profiles or strong assumptions about the behavior of these two profiles. In fact, it seems that legal experts require or, at least, expect this kind of comparison (FRA, 2018). [^7] We believe that this practice should not come as a surprise to the reader: if it takes more than one throw to check whether a coin is fair, why should it not be the same for a serious accusation like discrimination? Intuitively, we wish to control for chance from (or rule out out uncertainty in) the decision process in question to have (some) certainty on the pattern we are testing for. This is not possible under a literal comparison.
Together, these two definitions describe the general setup implicit to all (individual) discrimination tools. We provide an overview of these tools in Section 5. Given our focus on revisiting the standard comparator, in this work we focus on Definition 2.1. Clearly, whatever we derive as the comparator in (2) conditions what outcomes are compared in (3) and, in turn, what we test for and uncover as discrimination when using any discrimination testing tool.
2.2. On (Establishing) Discrimination
The comparator is intrinsic to how we conceive discrimination. This conception is comparative (Westen, 1982; Weerts et al., 2023; Kohler-Hausmann, 2019). We draw a distinction between describing versus proving something as discriminatory, which helps understand discrimination as a social phenomenon (the former) and as a modeling problem (the latter). Let us start with describing discrimination. Quoting from Weerts et al. (2023), who base their definition on Lippert-Rasmussen (2006): “discrimination can generally be characterized by the morally objectionable practice of subjecting a person (or group of persons) to a treatment in some social dimension that, for no good reason, is disadvantageous compared to the treatment awarded to other persons who are in a similar situation, but who belong to another socially salient group.” A group is considered to be socially salient “if perceived membership of it is important to the structure of social interactions across a wide range of social contexts” (Weerts et al., 2023). Defining the socially salient group, e.g., can only be done by agreeing on a specific, shared social context. Arriving to this point, e.g., by recognizing certain groups under non-discrimination law, captures the social phenomenon.
At the core of this view on discrimination there is a clear comparative element. We are not interested in the difference in treatment between any two individuals from different socially salient groups, but in the difference in treatment between two similar (or similarly situated) individuals from different socially salient groups. Hence, discrimination occurs when similar (or similarly situated) individuals that differ on membership to a socially salient group are treated differently. Under this setting, discrimination becomes the opposite of equality (Weerts et al., 2023). Also under this setting, the comparator becomes a manifestation on what similar (or similarly situated) individuals look like.
Let us now address proving discrimination. We consider the European Union (EU) context because, one, most fair ML works consider only the US context and, second, the EU—with the AI Act and GDPR—currently leads on the regulation front for algorithms (Álvarez et al., 2024). In doing so, we want to show the complex pipeline behind establishing discrimination, and emphasize that it does not begin nor end with Section 2.1. Beyond that, although the discussion below concerns EU non-discrimination law, the modeling problem formalized in Section 2.1 is shared among most (Western) conceptions of justice (Westen, 1982) and it is, thus, a robust problem to the legal context. There are two forms of discrimination conceived by EU non-discrimination law: direct and indirect discrimination. Under direct discrimination, the decision-maker uses the protected attribute for the decision, while under indirect discrimination the decision-maker uses a supposedly neutral attribute that is informative of the protected attribute. [^8] In US non-discrimination law, direct discrimination is refer to as disparate treatment and indirect discrimination as disparate impact (Barocas and Selbst, 2016). The EU context almost fully translates onto the US context, especially when it comes to direct discrimination and disparate treatment. For our purposes, the goal of substantive equality over formal equality is specific to the EU (Wachter et al., 2020). We come back to this point in Section 3, but, regarding the overall problem of the comparator, we can generalize across the two legal contexts. The same seems to hold for UK non-discrimination law (Sargeant and Magnusson, 2024). There are four elements in a discrimination claim, which we list below in order (Weerts et al., 2023). We add in italics the implications to modern decision flows.
  • “On grounds of”… We need to determine whether the claim falls under direct or indirect discrimination, which will depend on whether the decision is taken based on the protected attribute or not. This element has been interpreted in terms of whether the protected attribute is or not an input to the ML model (Hacker, 2018): currently, models are not allowed to use protected attributes. Recent works argue that if the model does not use the protected attribute, but is able to infer it (read, proxy discrimination (Tschantz, 2022)), then it should still fall under direct discrimination (Adams-Prassl et al., 2022).
  • …“a protected characteristic.” The claim must be based on an attribute protected by non-discrimination law. This element, on top of the current discussions on intersectional versus multiple discrimination (Xenidis, 2020), is ongoing as there is a fear that models are creating new protected attributes.
  • … where there is evidence for “less favorable treatment”… The complainant, to establish a case, needs to show prima facie evidence, meaning “sufficient evidence for a rebuttable presumption of discrimination to be established by the judge” (Weerts et al., 2023). If prima facie discrimination is established, the burden of proof falls on the defendant. Under a ML model, evidence itself becomes standardized. In principle, we must focus on the input-output pair of the model in question.
  • …unless there is an “objective justification.” Direct discrimination cannot be justified in principle. Indirect discrimination, instead, can be justified as long as it has a legitimate goal and passes the proportionality test. As Weerts et al. (2023) point out, neither the law provides concrete guidelines for the proportionality test nor can it be settled in advance. Further, the “objective justifications” are settled on a case-by-case basis. This element shows that we can train a ML model that could be used across multiple cases but to judge its fairness (for the purpose of testing discrimination) will depend on each case (Wachter et al., 2021).
The third element is the one that concerns the comparator as it requires the use of the discrimination testing tools for obtaining the prima facie evidence. Proving discrimination, similar to describing discrimination, has a clear comparative element; here, though, it has a modeling angle because the goal of obtaining evidence is explicit. The idea is still to compare similar (or similarly situated) individuals, though explicitly by defining $d$ in Definition 2.1, materializing what similarity between individuals is within our context.
Defining similarity is precisely the main challenge inherent to modeling the comparator $i^{\prime}$ with respect to the complainant $i$. The social context that describes discrimination can be and often is taken for granted when modeling the comparator as it is well understood. In Example 1.1, e.g., we did not need to motivate females and males as socially salient groups. The same occurs with most fair ML papers that simply list protected attributes like race and gender to which the proposed methods apply (Ruggieri et al., 2023). However, a well-understood context still does not release us from making normative statements that are socially charge when we define similarity between individuals through $d$. We cannot escape defining $d$ either as only through $d$ we derive the comparator $i^{\prime}$ for the complainant $i$. In principle, as long as some suitable $d$ exists, the modeling problem of proving discrimination appears simple and even intuitive. The simplicity, however, breaks down precisely because we must first define similarity between individuals to then obtain similar individuals and, in turn, prove discrimination. In The Empty Idea of Equality, Westen (1982) points out at this limitation of non-discrimination law by arguing that equality is a circular concept: to argue for equality we first must define what it means to be equal. It follows, thus, that similarity is a circular concept. The circularity of similarity matters because, although a discrimination tool’s main goal is to find evidence, which on paper might seem an objective goal, any discrimination testing tool—be it implicitly or explicitly—needs to implement a normative statement on similarity in the form of some $d$ in order to be operational. It is a difficulty inherent to deriving the comparator, be it the standard one or not. Hence, none of these tools can be sold as unbiased in practice.
Given that $d$ is never fully objective, a valid question for concluding this section is whether similarity can be ever defined without mathematical formalism? It is possible, but not the common approach, especially when it comes to testing discrimination. In philosophical works on discrimination (see, e.g., Loi et al. (2023)), the question of similarity is addressed through argumentation. The main drawback we see with such an approach is the lack of scalability. Granted that mathematical formalism may give a false sense of objectivism relative to argumentation (Kohler-Hausmann, 2019; Hu and Kohler-Hausmann, 2020), the fact remains that establishing discrimination requires evidence. Simply put, if we need to find or generate other similar individuals to the complainant in order to build our evidence, then using $d$ is more convenient than arguing the similarity of multiple individuals to the complainant on a case-by-case basis.
2.3. The Counterfactual Model of Discrimination
Discrimination is often conceived as a causal claim on the (in)direct effect of the protected attribute $A$ on the decision outcome of interest $Y$ (Heckman, 1998). In Example 1.1, e.g., we are interested in whether Clara did not receive tenure because of her gender? Its causal underpinning, which is motivated by non-discrimination law’s own definition of discrimination (Weerts et al., 2023), has long motivated a range of methods for testing discrimination based on counterfactual reasoning. These methods operationalize the scenario in which we are able to manipulate the protected attribute of the individual(s) making the claim, imagine the “what would have been if” (or the counterfactual) outcome, and compare it to the “what is” (or the factual) outcome to isolate the causal effect of $A$ on $Y$. Kohler-Hausmann (2019) refers to this practice as the counterfactual model of discrimination. [^9] Kohler-Hausmann (2019) actually refers to it as the counterfactual causal models of discrimination, which is redundant as counterfactuals are causal (Woodward, 2005). Hence, we drop the “causal.” Using SCM, we can state the discrimination claim as $A\rightarrow Y$ (if direct) or as $A\rightarrow\mathbf{X}\rightarrow Y$ (if indirect). The goal of any discrimination testing tool is validating such claim.
What we want to emphasize about the counterfactual model of discrimination is not that it relies on counterfactual reasoning to answer causal claims, but that, because of this reliance, the comparator $i^{\prime}$ is a counterfactual representation of the complainant $i$. Viewing $i^{\prime}$ as a counterfactual representation of the factual $i$ ultimately positions the comparator as a causal modeling problem. This additional emphasize from our side is important when we consider that $i^{\prime}$ can be either a real or fictitious profile, making it hypothetical in the fullest sense. That the comparator $i^{\prime}$ can be viewed as a thought experiment is not exploited–both conceptually and experimentally–under the CP comparator. We come back to this point in Section 3. We diverge from Kohler-Hausmann (2019) precisely because of this emphasize, viewing the dominant counterfactual model of discrimination as an opportunity suited for fair ML, not as a threat.
Why counterfactuals? We argue, one, because discrimination represents a pattern and, two, because the discrimination claim requires an answer at the individual level. Let us unpack each point. First, why would discrimination by being a pattern amount to a causal relation? Broadly, the presence of a sufficiently robust pattern suggests a causal relationship, as causality is seen as an invariant force acting on the world. [^10] Formally, such view amounts to an interventionist account of causality, which is the basis for both SCM Pearl (2009) and Potential Outcomes Angrist and Pischke (2008), meaning the two most popular causal modeling frameworks. See Woodward (2005) for details. The presence of a causal pattern also suggests, implicitly, that there is little room for chance to explain what we observe. Discrimination claims are often characterized in those terms, meaning an unfair decision that persists over time toward a certain group of individuals. Because of this persistence, we are (somewhat) certain that the unfair treatment comes from the influence of the protected attribute. Intuitively, a discriminatory decision-making process has (un)consciously set up a pattern (i.e., $A\rightarrow Y$ or $A\rightarrow\mathbf{X}\rightarrow Y$) that causes the unfair treatment of protected individuals. If the goal is to detect such pattern(s) and we agree on the premise that discrimination represents a causal pattern, then counterfactual reasoning is the main approach to do so. There is a long tradition in causality, especially within SCM, to formulate what constitutes a cause in terms of counterfactual reasoning (Pearl, 2009; Halpern, 2016; Halpern and Pearl, 2001a, b): if intervening $A$ induces a counterfactual outcome $Y^{CF}$ different from the factual outcome $Y$, then $A$ must be a cause of $Y$ [^11] These definitions are much more formal, such as, in terms of probability distributions, but the phrase captures the main idea behind all of them. Under this setting, to test for the existence of such pattern would imply to examine the impact of $A$ on the decision-making process by changing $A$ itself. If there is such a pattern, then observing a change in outcome due to having changed only $A$ (and controlling for everything else) would, first, confirm the existence of the pattern, and, second, confirm the role of $A$ within it.
Second, why would counterfactual reasoning and, in turn, the comparator, answer individual-level queries? Although it is trivial to point out, no individual $i^{\prime}$ can be more similar to another individual $i$ than that same individual $i$. When testing for discrimination we are trying to imagine how that same individual would have been like in another life. This is the sort of mental exercise we carry out when asking “what would have been of Clara had she been male?” in Example 1.1. It is a hypothetical question that, regardless, we aim and, to an extent, need to answer with discrimination testing. We cannot observe simultaneously, for any individual $i$, both his or her “what is,” or factual, and “what would have been if,” or counterfactual, outcomes. We are only certain of the former, factual outcome; and the best we can do is to derive another individual $i^{\prime}$ that represents (or approximates) the latter, counterfactual outcome of $i$. Defining similarity between individuals through $d$ to obtain such $i^{\prime}$ is an attempt at controlling for all factors that might influence an outcome and being able to answer whether one of those factors, i.e., $A$, is a cause of $Y$. We stress that this line of reasoning is implicit to all causal and non-causal discrimination testing tools using the counterfactual model of discrimination (Kohler-Hausmann, 2019). Under the counterfactual model of discrimination, similarity is not only a statement between two individuals $i$ and $i^{\prime}$, but it is also a statement on the alternative paths attributed to an individual $i$ through the lived experiences of another (real or fictitious) individual $i^{\prime}$. Similarity, in turn, becomes also a statement on this factual world and some hypothetical counterfactual alternative.
3. Mutatis Mutandis
In this section, we revisit the CP comparator of Definition 2.1, which is the standard in discrimination testing, by introducing the mutatis mutandis, or “with the appropriate adjustments being made,” (MM) comparator. We first define the MM comparator below. We then motivate it and discuss how it extends the problem of discrimination testing.
Definition 3.1.
(The MM Comparator) Let $\tilde{d}$ be a protected-aware distance function that, unlike $d$ in (2), considers both $\mathbf{X}$ and $A$ as inputs, capturing the potential influence of $A$ on $\mathbf{X}$. For the complainant $i$ $\langle\mathbf{x}_{i},a_{i},y_{i}\rangle$, we define a mutatis mutandis (MM) comparator $i^{\prime}$ as an individual profile $\langle\mathbf{x}_{i^{\prime}},a_{i^{\prime}},y_{i^{\prime}}\rangle$ such that:
(4) $\tilde{d}((\mathbf{x}_{i},a_{i}),(\mathbf{x}_{i^{\prime}},a_{i^{\prime}}))\leq\epsilon\;\;\;\text{and}\;\;\;a_{i}\neq a_{i^{\prime}}$
where ϵ+italic-$\epsilon\in\mathbb{R}^{+}$ is a threshold.
Under $\tilde{d}$, two instances with different non-protected attributes, $\mathbf{x}_{i}\neq\mathbf{x}_{i^{\prime}}$, can be considered similar given information, respectively, about $a_{i}$ and $a_{i^{\prime}}$. The fact that Definition 3.1 allows for this “dissimilarity” between comparator and complainant motivates the MM denomination. The general idea is that under $\tilde{d}$ we should be able, if needed, to account for the influence of $A$ on $\mathbf{X}$ when deriving the $i^{\prime}$ relative to $i$. In other words, we “adjust for what needs to be adjusted” as a consequences on “changing” $A$, allowing us to derive an individual profile less conservative than the one conceived under $d$ in Definition 2.1. We motivate this argument in the next sub-sections.
Given the meaning behind $\tilde{d}$, note that, when $a_{i}=a_{i^{\prime}}$, the MM and CP comparators are the same. As we move within the same group in $A$, we assume that the influence of $A=a$ on $\mathbf{X}$ is shared among all $i$ individuals with $A=a_{i}$. Implicitly, we assume that defining similarity between individuals of the same group in $A$ requires no additional considerations due to shared experience. Additionally, when $\mathbf{X}$ and $A$ are independent, the MM and CP comparators are the same even if $a_{i}\neq a_{a^{\prime}}$ as there are no downstream effects of $A$ to $\mathbf{X}$ to be accounted for.
3.1. Against Idealized Comparisons
Let us focus on the standard, CP comparator. We argued already that any notion of similarity between individuals implemented through the distance function $d$ in (2) is a normative statement. We also argued that, in turn, the comparator derived under $d$ represents the normative statement. This is true for the standard comparator as for any other formulation of the comparator, such as the MM comparator. Now there are two characteristics specific to the CP comparator worth stressing. First, in Definition 2.1, similarity between individuals comes down to a comparison on the non-protected attributes $\mathbf{X}$ between $i^{\prime}$ and $i$. The distance function $d$ is oblivious to (or unaware of) the protected attribute $A$. This holds whether $A$ and $\mathbf{X}$ are associated or not as $d$ does not take it into consideration. Second, it follows that the counterfactual representation of the complainant $i$ in the form of the comparator $i^{\prime}$ only varies, ideally, in terms of $A$. In other words, nothing would have changed for $i$ in $\mathbf{X}$ had she or he had been from the non-protected group. Because of these two characteristics, we say that the standard CP comparator represents an idealized comparison between $i^{\prime}$ and $i$.
Is the idealized comparison problematic, though? The answer to this questions comes down to having a stance on similarity between individuals, and how it should be operationalized through the $d$ in Definition 2.1. We view it as open to discussion. In this work, we do not have a strong stance against the CP comparator as it can be useful, e.g., in uncovering direct discrimination or when an alternative comparator is not possible. We do, however, view the discussion against it as necessary for discrimination testing. This discussion, e.g., has motivated the MM comparator.
Key in criticizing the CP comparator and the idealized comparison it represents is Kohler-Hausmann (2019). Under the premise that the comparator $i^{\prime}$ in Definition 2.1 represents a counterfactual representation of the complainant $i$, Kohler-Hausmann (2019) criticizes the counterfactual model of discrimination for ignoring how membership to a protected group conditions the material outcomes of individuals and, thus, limits any meaningful counterfactual analysis between protected and non-protected individuals. Kohler-Hausmann (2019), in particular, argues that protected attributes, like race and gender, are social constructs, which is a view widely held by social scientists (Bonilla-Silva, 1997; Sen and Wasow, 2016; Hanna et al., 2020). What it means is that these social categories were once (or still are) used not only to classify individuals but to also delimit the opportunities available to them (Mallon, 2007). According to Kohler-Hausmann (2019), it is inconceivable to picture a counterfactual white version of a black individual as stated in Definition 2.1. Under such definition, the ideal comparator $i^{\prime}$ will have a set of non-protected attributes identical to those of $i$, or $d(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})=0$. By envisioning a counterfactual world where $i$ can be exactly as he or she is today, in terms of non-protected attributes, while belonging to the non-protect group, this mental exercise requires from us to accept that the protected individual $i$ would have arrived at the same point had he or she been a non-protected individual $i^{\prime}$, which reduces the pervasive influence of $A$.
Following Kohler-Hausmann (2019)’s constructivist argument, Definition 2.1 envisions a scenario where we can manipulate $A$ and expect $\mathbf{X}$ to remain unchanged, failing to capture the role of $A$ and raising the question around what exactly are the discrimination testing tools assessing for. As she writes in the case of race:
The problem with identifying discrimination with the treatment effect of race is that it misrepresents what race is and how it produces effects in the world, and concomitantly, what makes discrimination of race a moral wrong. […] [I]f the signifiers of racial categories fundamentally structure the interpretation and relevance of other characteristics or traits of the unit, then it is a mistake to talk about identical units that differ only by raced statues.
In short, Kohler-Hausmann (2019) implies that the comparator $i^{\prime}$ derived in Definition 2.1 is wrong and, in turn, the tools for testing discrimination built using such comparator are also wrong. [^12] Kohler-Hausmann (2019) later extends the above argument for gender (Hu and Kohler-Hausmann, 2020), covering the two main protected attributes. We agree partially with her. First, we do not see all discrimination testing that uses the CP comparator as wrong as it might be the only viable options in many cases. Second, we do not picture a better modeling framework for discrimination testing than the counterfactual model of discrimination. It is not ideal, but it can also improve if we are willing to revisit the derivation of the comparator.
Another critical work toward discrimination testing worth mentioning is Michaels and Reed’s article, The Trouble with Disparity (Michaels and Reed, 2020). We believe that it offers a similar critique to Kohler-Hausmann (2019), but from a non-legal angle. Michaels and Reed, in particular, focus on the tension between race and class and use the coverage given to black communities during the COVID-19 pandemic to illustrate their points. During the pandemic in the USA, blacks were dying from the virus at a higher rate than any other social group, which prompted “scientific” questions on whether there was a genetic pre-disposition that made this group more vulnerable to the virus relative to other groups. No such evidence was found. Once we controlled for other factors to isolate the so-called treatment effect of race, it became clear that “race” had nothing to with the likelihood of dying from the virus. Drawing parallels to non-discrimination law and its focus on parity based on representation, Michaels and Reed (2020) point at the obvious or, in their words, deeper cause that nobody wanted to address regarding the disparate death rates among social groups from COVID-19: inequality. Blacks were not more likely to die from COVID-19 because of the color of their skin but because of their socioeconomic background. Being poor is what made an individual in the USA more vulnerable to COVID-19, and the majority of poor individuals in the USA are black.
We mention this article because it highlights the complexity of discrimination testing. The answer to “why are most low-income individuals in the USA today black?” is the same answer to “why is race considered a protected attribute?”: because there is a history of systematic policies against individuals perceived as black (or, in general, non-white). Now, because of that same answer, the question we aim to answer when testing for the discriminatory effects of race should not be limited to race alone. The same way we have established that race and other protected attributes exist and have acknowledge their effects in our society, we must also recognize how their effects have materialized in limiting the opportunities of multiple generations. Back to the COVID-19 example, it might have been race what initially caused the exclusionary policies toward non-whites that, in turn, created the current social context, but, precisely because of this historical process, race alone cannot claim full responsibility of the present disparities. Larger forces, though, still associated to race, such as inequality and social mobility, are stronger causes to the present disparities. See, e.g., Chetty et al. (2020). The problem is that, ironically, as Michaels and Reed argue, it has become easier to speak of racial discrimination than income discrimination.
To be clear, arguments against discrimination testing are not new. For instance, Heckman (1998) already criticized ideal comparisons in causal modeling. Similarly, Wachter et al. (2020), by classifying fair ML methods as bias preserving or bias transforming depending on how they define similarity, recently pointed out at the limitations of using the conservative CP comparator. These arguments, however, are largely absent from the discussion around discrimination testing tools. There, the focus is always on tool development, which is to be expected. However, as we already argued, the circularity of similarity requires for a deeper considerations from the tool developers.
3.2. A New Kind of Comparator
Once we allow for a more complex view on the protected attribute $A$ and its influence on the non-protected attributes $\mathbf{X}$, by, e.g., viewing $A$ as a social construct (Kohler-Hausmann, 2019) or acknowledging $A$ as a source of inequality (Michaels and Reed, 2020), the CP comparator appears insufficient. Under this view, there is the need to clarify the relationship between $A$ and $\mathbf{X}$, and whether it is in our interest to account for it when testing for discrimination. The MM comparator, based on $\tilde{d}$ in (4), is meant to clarify and account for the relationship between $A$ and $\mathbf{X}$, embodying what is refereed to as “fairness given the difference” (Álvarez and Ruggieri, 2023). We believe that the MM comparator represents a new kind of comparator, aimed at not only proving discrimination but at shaping the social context behind it.
The MM comparator represents an ideal target level of equality instead of the formal one captured by the CP comparator due its idealized comparison. Conceptually, if $\tilde{d}$ accounts for the influence of $A$ on $\mathbf{X}$, then by comparing $i^{\prime}$ to $i$ when $\mathbf{x}_{i^{\prime}}\neq\mathbf{x}_{i}$ and $a_{i^{\prime}}\neq a_{i}$ we are making a normative statement claiming that these two observably dissimilar individuals should be similar. Under counterfactual reasoning, since the counterfactual question is around $A$, the comparator $i^{\prime}$ derived under Definition 3.1 represents where the complainant would have been if free from the effect of being part of the protected group. The CP comparator does not allow for this possibility as it only envisions a comparator $i^{\prime}$ all equal but for $A$ to the complainant $i$. That is why in Example 1.1, Mike is the CP comparator while Vincent the MM comparator, as Vincent, with his higher number of publications, is meant to represent a Clara that would have carried out her academic career without the social expectations of being the main caregiver of her household.
The MM comparator acts as a target because we, essentially, need to generate it. In Example 1.1, e.g., why should we consider Vincent when other profiles could also meet the MM comparator definition? In other words, we need to have a modeling procedure that motivates the choice of the MM comparator. In practice it means estimating the effect of $A$ on $\mathbf{X}$, and adjusting the latter accordingly. We discuss this last point in the next section. Under this “modeling” interpretation, the proposed MM comparator might seem abstract. In fact, why complicate the derivation of the comparator under the MM definition when we already have the CP definition? A reason is that, as argued by Wachter et al. (2020), non-discrimination law, at least in the EU , aims at substantive equality. In practical terms, it means that when testing for discrimination, we are not interested in whether the treatment was the same for similar individuals, but whether the setting (Wachter et al. (2020) refers to it as the “status quou”) is suitable for similar individuals to be treated similarly over time. This point is similar to the argument on why equal opportunity is better than demographic parity by Hardt et al. (2016): there is a difference between hiring equal number of males and female from ensuring that those numbers remain over time. In fact, this distinction summarizes why Wachter et al. (2020) classifies fair ML definitions into bias preserving (like demographic parity) and bias transforming (like equal opportunity).
A second reason to consider the MM comparator is that, implicitly, we are already implementing methods that aim at ideal targets. We are not proposing a drastic turn in fair ML research. A good example of this point is the fairness-accuracy trade-off (Wick et al., 2019; Sharma et al., 2020; Rodolfa et al., 2021; Maity et al., 2020; Dutta et al., 2020). Recent works on this topic agree that the trade-off exists as long as the data is biased. Hence, demonstrating its existence is trivial given the current state of our society; what is of interest is studying when or under which conditions the trade-off disappears. These are hypothetical conditions that, though, serve as future targets to aim for. Not surprisingly, in a setting in which $\mathbf{X}$ is non-informative of $A$ or, similarly, $A$ does not influence $\mathbf{X}$, there is no need to consider a trade-off between accuracy and fairness. In a fair world, the most accurate model is also the most fair. We believe that the MM comparator is meant to capture such world for the purpose of discrimination testing.
3.3. Extending the Setup
Consider indirect discrimination cases, meaning $A\rightarrow\mathbf{X}\rightarrow Y$ or $Y=f(\mathbf{X})$. Under the causal modeling problem of deriving a comparator, the key question to be answered is: what happens to $\mathbf{X}$ once we manipulate $A$ to derive the comparator? Given how Definitions 2.1 and 3.1 differ, assuming a case of indirect discrimination, the CP comparator ignores the downstream effect of $A$ on $\mathbf{X}$ while the MM comparator accounts for it. However, while the implementation of the former is clear, how do we actually implement the latter? How do we define $\tilde{d}$ in (4)? Answering this question helps extend the discrimination testing setup in Section 2.1. To start, let $g$ represent a generative model such that:
(5) $\text{for}\;\;\tilde{X}=g(\mathbf{X},A)\;\;\;\text{we have}\;\;\;\tilde{X}\perp\!\!\!\!\perp A$
We can view, broadly, $\tilde{X}$ as an adjusted representation of $X$ that is removed from the influence of $A$. Conceptually, (5) describes the problem of fair representation learning (Zemel et al., 2013) in which we want to learn or have access to $g$ in order to achieve $\tilde{X}$. Such a model, e.g., can be a Variational Auto-Encoder or a Markov Chain (Kingma and Welling, 2014; Salimans et al., 2015; Bengio et al., 2013).
For our purposes, we view $\tilde{X}$ as a counterfactual (fair) representation of $\mathbf{X}$ that answers to the question of what would have happened had the individual belonged to the non-protected group instead of the protected group? Hence, $g$ represents the counterfactual generation procedure described in Section 1.1. Given $g$, we should be able to generate the set of non-protected attributes of the complainant $i$’s comparator $i^{\prime}$ and derive the MM comparator $i^{\prime}$ such that:
(6) $d(\tilde{\mathbf{x}}_{i},\mathbf{x}_{i^{\prime}})\leq\epsilon$
where, possibly, $\tilde{\mathbf{x}}_{i}\neq\mathbf{x}_{i}$. Given the causal relation $A\rightarrow\mathbf{X}$, the comparator $i^{\prime}$ is similar to the counterfactual profile of $i$ but dissimilar to its factual profile. Here, $d$ represents again a distance function unaware of the protected attribute $A$ in (2). Through (6), we summarize the link between the Definitions 2.1 and 3.1 complainant $i$ and its comparator $i^{\prime}$ as:
(7) $\tilde{d}\big{(}(\mathbf{x}_{i},a_{i}),(\mathbf{x}_{i^{\prime}},a_{i^{\prime}})\big{)}=d\big{(}g(\mathbf{x}_{i},a_{i}),\mathbf{x}_{i^{\prime}}\big{)}$
where the left-hand-side of the equation represents the general definition of the MM comparator while the right-hand-side an implementation of it under the generative model $g$. The relationship (7) allows to derive the MM comparator via $\tilde{d}$ or via $g$ plus $d$. Our preference, based on our counterfactual reasoning approach, is with using $g$ plus $d$, which allows for: (i) making it clear the difference between the MM and the CP approaches; (ii) reusing the existing notions of distances and algorithms for inferring them (Kulis, 2013; Ilvento, 2020). In that case, the setup in Section 2.1 would require an additional element on top of deriving the comparator and comparing the outcomes: learning $g$, which is of course not limited to counterfactual generation. This extended setup for discrimination testing represents an opportunity for ML generative models and a promising path for future work.
4. An Illustrative Experiment
[…]
$\mathcal{M}\,$$\begin{cases}R&\leftarrow U_{R}\\ G&\leftarrow U_{G}\\ UGPA&\leftarrow b_{U}+\beta_{1}\cdot R+\lambda_{1}\cdot G+U_{1},\\ LSAT&\leftarrow\exp\{b_{L}+\beta_{2}\cdot R+\lambda_{2}\cdot G+U_{2}\}\\ \end{cases}$
$Y$ =$=\begin{cases}1\;\text{if}\;(0.6\cdot UGPA+0.4\cdot LSAT)>\psi\\ 0\;\text{otherwise}\\ \end{cases}$
Figure 1. The causal knowledge with corresponding SCM $\mathcal{M}$ and DAG $\mathcal{G}$ for Law School Admissions. Let $b_{U}$ and $b_{L}$ denote the intercepts; $\beta_{1}$, $\beta_{2}$, $\lambda_{1}$, $\lambda_{2}$ the weights; and $UGPA\sim\mathcal{N}$ and $LSAT\sim\text{Poisson}$ the probability distributions.
To showcase the CP (Definition 2.1) and MM (Definition 3.1) comparators and their impact on discrimination testing, we consider the Law School Admissions example popularized by Kusner et al. (2017) and use on it the discrimination tools of situation testing (ST) (Thanh et al., 2011) and counterfactual situation testing (CST) (Álvarez and Ruggieri, 2023). ST and CST are data mining tools that look for $k$-nearest neighborhoods for complainant-comparator pairs in a dataset. We focus on these two tools because CST extends ST by implementing the MM comparator over the CP comparator. Both tools follow a similar pipeline, even sharing the same distance function for constructing the neighborhoods around the complainant and comparator, with the key difference that ST flips the protected attribute to create its (CP) comparator, whilst CST uses the generated counterfactual of a complainant $i$ to create its (MM) comparator $i^{\prime}$. The code developed for this section is publicly available in a Github repository. [^13] Visit: https://github.com/cc-jalvarez/revisiting-the-comparator.
The dataset contains $n=21790$ applicants, containing the applicants’ gender ($G$), race ($R$), undergraduate grade-point average ($UGPA$), and law school admissions test score ($LSAT$). In terms of the protected attributes, 43.8% of applicants are females and 16.1% are non-whites. Following Álvarez and Ruggieri (2023), we define a decision-maker $b$ that uses only $UGPA$ and $LSAT$ for admitting $Y=1$ or not $Y=0$ an applicant. The decision-maker $b$ test if a weighted sum of $UGPA$ and $LSAT$ is above a cutoff $\psi$, where $\psi$ is the median entry requirement of a top US law school.
We are interested in testing, separately, whether female and non-white applicants are potentially discriminated under this decision-making process. [^14] Intersectionality, meaning considering what occurs with non-white-female applicants, is not considered under most non-discrimination laws (Xenidis, 2020; Crenshaw, 1989). Testing-wise, it means that we can only consider the claims for gender and race separately. To illustrate these tools, consider a female applicant as the complainant $i$. ST finds $k$-females and $k$-males with the closest $UGPA$ and $LSAT$ to $i$ and compare the difference in the proportion of rejections between both groups. The complainant is considered discriminated if such a difference is significantly different from zero at $95\%$ confidence level. Under ST, the ideal comparator $i^{\prime}$ in ST is a male with the same $UGPA$ and $LSAT$ as $i$. CST does the same with the exception that it first generates $i$’s counterfactual using Figure 1 and uses it to find the $k$ males $UGPA$ and $LSAT$. The female group is construct the same way as in ST as the counterfactual generation does not apply to them. Under CST, there is no ideal comparator $i^{\prime}$ to be found but, instead, a counterfactual representation of $i$ in the form of $i^{\prime}$ with $\tilde{\mathbf{X}}$ (6). Because of how each method constructs the male neighborhoods, ST and CST implement, respectively, the CP and MM comparators.
Table 1. Number (and % over the total cases) of discrimination cases based on race.
Method $k=25$ $k=50$ $k=100$ $k=200$ $k=500$
ST 40 (1.14%) 61 (1.74%) 64 (1.83%) 75 (2.14%) 108 (3.08%)
CST 296 (8.44%) 337 (9.61%) 400 (11.41%) 476 (13.58%) 605 (17.26%)
Table 2. Number (and % over the total cases) of discrimination cases based on gender.
Method $k=25$ $k=50$ $k=100$ $k=200$ $k=500$
ST 117 (1.23%) 229 (2.4%) 258 (2.71%) 368 (3.86%) 492 (5.16%)
CST 106 (1.11%) 253 (2.65%) 296 (3.1%) 449 (4.71%) 605 (6.34 %)
Tables 1 and 2 show the number of discrimination cases identified under ST and CST for, respectively, the protected attributes race (what would have been the outcome had the applicant been white?) and gender (what would have been the outcome had the applicant been male?) for different $k$ neighborhood sizes. [^15] The CST results are without the complainant-comparator pair. This is because ST does not include them, which allows to compare the results for both methods. Results including the pair are similar to those without it. See Álvarez and Ruggieri (2023) for details. Both tables illustrate the impact of choosing between the CP and MM comparators, with CST detecting a higher number of discrimination cases for the same dataset. This difference is driven by ST, under the CP comparator, and CST, under the MM comparator, aiming, respectively for an idealized and a fairness given the difference comparison. Intuitively, under a stricter or more conservative notion of similarity, the ST constructs similar neighborhoods based on the protected attribute. We observe this in Figure 2 that shows seven randomly chosen individuals considered to have been discriminated by both ST and CST on the basis of their gender. We compare the $LSAT$ and $UGPA$ distributions in the “test groups” (i.e., the neighborhoods constructed around the comparator) of ST and CST to the $LSAT$ and $UGPA$’s distribution in the “control group” (i.e., the neighborhood constructed around the complainant), which is the same for both ST and CST. The two groups under ST for $LSAT$ and $UGPA$ are similarly distributed while this is not the case under CST. Figure 2 illustrates $\tilde{\mathbf{X}}$ (6) and its role in discrimination testing. In Figure 3 we carry out the same analysis and find similar results for race.
Under CST, the complainant used for drawing the so-called test groups is purely generated under counterfactual generation. In other words, these profiles determining which male (or white) outcomes are to be compared to the female (or non-white) outcomes are purely hypothetical. Each of these generated profiles represents what would have been of the female (or non-white) individual had it been male (or white) in the admissions process to law school. In generating them and, in turn, adjusting for the impact of gender (or race) on $UGPA$ and $LSAT$, CST creates individual targets for each protected individual that represent where they would be in a fairer (read, free from the impact of the protected attribute) world. CST, thus, tests for equality under these targets. Of course, these results are specific to ST and CST and the potential limitations inherent to both methods, though they still highlight what it would mean to test for discrimination under the MM comparator instead of the CP comparator.
Figure 2. We show the distributions in the form of box-plots for the $LSAT$ and $UGPA$ for seven randomly chosen complainants based on the protected attribute gender. These complainants are found to be discriminated by both ST and CST for $k=100$. The so-called control (ctr) are the neighborhoods of similar female profiles to the complainant, which is the same for both tools. Similarly, the so-called test groups (tst-st for ST; tst-cf for CST) are the neighborhoods of similar male profiles to the comparator: i.e., the CP comparator for ST, and the MM comparator for CST. In both figures, we observe a difference between the two kinds of test groups. As we would expect, it means that the test and control groups are more similar under ST than under CST.
Figure 3. Here, the same analysis follows as in Figure 2 but for the protected attribute race, including similar insights and conclusions on the influence of the CP and MM comparators. In both figures, we observe a difference between the two kinds of test groups.
5. Overview of Discrimination Testing Tools
In this section, we survey a representative set of tools for discrimination testing, focusing on the comparator. We recommend Romei and Ruggieri (2014) for an extensive and multidisciplinary survey. All these tools amount to gathering prima facie evidence against a decision-maker, meaning they derive the comparator $i^{\prime}$ for the complainant $i$ and compare their outcomes. We view these tools as instances of the counterfactual model of discrimination.
5.1. Standard Tools
By standard we mean non-algorithmic tools. These tools mainly include natural experiments (e.g., Goldin and Rouse (2000)), field experiments (e.g., Bertrand and Duflo (2017)), and audits (e.g., Fix and Struyk (1993)). These tools share a focus on either gathering or evaluating data that is, by design, able to capture the effect of the protected attribute on the decision outcome. These tools are somewhat outdated for how decisions are taken in modern decision flow.
Let us start with audits, which are a more holistic kind of tool, very human-dependent, and less focused on modeling. Audits consist in examining the decision process in question through human auditors. In practice, it means sending a group of experts to observe during a period of time and study the practices of the decision-maker to corroborate the claims of the complainant (Fix and Struyk, 1993). In the case of the modern decision flow, such a practice only makes sense if the model is used by a human decision-maker or if the focus is to audit the procedure by which the model was trained (Koshiyama et al., 2024), which is increasingly being viewed as a priority over interpretability by regulators (Panigutti et al., 2023).
Natural and field experiments, instead, tend to have more of a focus on discrimination as a modeling problem (i.e., identifying the causal effect of the protected attribute) and, thus, prioritize designing and gathering data suitable for this purpose. These methods are a product of their time. In particular, these methods are a product of the so-called “empirical revolution” experienced by the social sciences in which a mixture of data availability, better computers, and an intuitive causal framework (the Potential Outcomes Framework) made it possible to test causal claims via experiments (Angrist and Pischke, 2008).
The overall idea behind natural experiments is simple. A policy introduction (or external shock) occurs creating a before-and-after scenario that, in principle, allows us to test for the effect of the protected attribute and find ourselves, naturally, in an experimental setting. Goldin and Rouse (2000)’s Orchestrating Impartiality: The Impact of“Blind” Auditions on Female Musicians is a clear example. US orchestras before the 1970s used to audition musicians without screens, meaning musicians would perform in full sight to the evaluating committee. The US orchestra system back then were male dominated and, some, openly sexist, which motivated the suspicion that there was discrimination against female musicians. Orchestras then started to conduct blind auditions to ensure an impartial process. Some orchestras even introduced carpets so that female musicians would not be identified by the sound of their heels when walking up the stage. Female representation in orchestras increased after the introduction of the screen. Goldin and Rouse (2000), though, using data from the auditions, set out to study whether this amounted to evidence for discrimination within the US orchestra system. There was indeed a before-and-after mechanism through the introduction of the screen that coincided with an increase in female musicians in orchestras, but to claim a causal effect we needed to control for other factors. For instance, the introduction of the screen also coincided with an overall increase of females in the US work force. Goldin and Rouse (2000) used a regression model to control for this and other factors, finding inconclusive evidence for discrimination against female musicians. Goldin and Rouse (2000) is a seminal paper, among other things, because it shows how meticulous the model choice has to be to claim discrimination evidence. It illustrates, implicitly, the importance of deriving a comparator good-enough to isolate the effect of the protected attribute.
Field experiments, instead, focus on designing an experiment to create the proper data for testing the discrimination claim. The most famous example is Bertrand and Mullainathan (2004)’s Are Emily and Greg More Employable Than Lakisha and Jamal? A Field Experiment on Labor Market Discrimination. Also see Bertrand and Duflo (2017) for a survey on field experiments for discrimination. Bertrand and Mullainathan (2004) proposed the following experiment: let us create pairs of fictitious CVs tailored to job positions that are the same except for the candidate’s name, where the name is intended to reflect both gender and race within the US context. They would create, e.g., the same CV for Mike and Jamal (respectively, a common white and black name according to census data at the time); send both CVs to the same job openings via mail (this was done in the early 2000s); and compare which CV received a callback. The CV with the black-sounding name was the complainant and the CV with the white-sounding name its comparator. All the CVs were drafted by the researchers. In the end, they found evidence for discrimination. This sort of field experiment is known as a correspondence study. See Rooth (2021) for a survey on correspondence studies. Another kind of field experiment is situation testing (Rorive, 2009) in which, rather than sending CVs or other forms of correspondence, we send a group of actors that share key physical traits and backstories as the complainant and its comparator.
Standard tools are at the center of the criticism against the counterfactual model of discrimination. This fact is not surprising, as these tools have been around longer, even inspiring the recent algorithmic tools. Heckman (1998), e.g., is critical of the effectiveness of these methods in controlling for similarity and other factors and argues that we have been too confident in claiming to be able detect discrimination. More recently, Kohler-Hausmann (2019) goes after these methods (she is particularly critical of Bertrand and Mullainathan (2004)) due to their experimental designs. These standard tools rely on the CP comparator. As previously discussed, under direct discrimination, these tools are still useful and have a lot to offer to algorithmic tools in terms of their focus on inferential statistics. If, however, we are interested in proving indirect discrimination, then these experiments have to make a strong argument for using the CP comparator. In particular, this argument should be centered on the kind of signal (as in relevant information) being used by the decision-maker and whether that signal has any link to the protected attribute. Consider the case of the orchestra auditions. There, the decision-maker cares only for the playing quality of the musician, which should not be related to gender. [^16] Though there is a reasonable case to be made here that instruments tend to favor male musicians (Criado-Perez, 2019). Hence, the signal can be recovered once the screen is imposed as the sounds can go through the screen but all other factors, like the musician’s appearance, are blocked. Consider now the case of the CVs. There, even if we impose a “screen” by removing the names, there is still the possibility that the decision-maker uses signals for candidate potential that are also linked to race or gender, such as extracurricular activities or the name of the high-school. These are valid signals, or at least signals that could be well argued by the decision-maker as necessary for measuring candidate potential. In such a setting, the MM comparator might be better suited.
5.2. Discrimination Discovery
Discrimination discovery tools (Kamiran et al., 2013; Ruggieri et al., 2010; Pedreschi et al., 2008) represent the first attempt at what we now call algorithmic fairness. These works took the standard methods and combined them with data mining methods, shifting the focus from generating or finding the experimental data to exploring a data produced by a decision process. In part, this shift in focus was field-driven: within data mining and overall knowledge discovery, the focus of the problem formulation is on developing new and efficient algorithms for extracting information from large datasets. This is a different focus from the prevalent inferential statics approach in which the creation of the data is as important as using the most efficient model. It came with a change in how we test for discrimination, especially with the adoption of modern decision flows, as the attention was now on a consistent, algorithmic decision-maker. Hence, to test whether the decision-maker discriminated, we could now look at the data it produced.
Under these tools, the setup in Section 2.1 started to become standard. It was a natural transition. Both the decision-maker in question and the tools for evaluating it were scalable. These works borrowed considerably both from non-discrimination law and from the standard methods. Works by Pedreschi et al. (2008) and Ruggieri et al. (2010) provided the first formalization to equal opportunity (and, implicitly, to equalized odds) considerably before Hardt et al. (2016). Kamiran and Calders (2009) did the same for demographic parity. Thanh et al. (2011) is an illustrative example of the discrimination discovery wave, proposing a $k$-NN implementation of situation testing.
These tools focus on defining the best distance to reach the comparator for a given complainant. Some of them are closer to the MM comparator than to the CP comparator, though. Thanh et al. (2011) clearly use the CP comparator as they minimize the distance between the complainant and potential comparators based only on the set of non-protected attributes. Zhang et al. (2016), instead, offer an extension to (Thanh et al., 2011) via causal weights, focusing on defining a distance function that prioritizes the important non-protected attributes when finding the comparator for a complainant. Similarly, Qureshi et al. (2020) use a propensity score weight for the distance function with the protected attribute as the treatment. None of these tools, however, explicitly address the problem with using the CP comparator or the need to derive the MM comparator. There is no discussion on how the protected attribute might influence all other attributes.
5.3. Algorithmic Fairness
In the past decade, discrimination discovery took a broader, more multidisciplinary focus and became algorithmic fairness. The objectives are the same, but algorithmic fairness is broader in scope, usually aiming to reformulate standard ML problems under a fairness objective (Barocas et al., 2019). In turn, many of these tools, even if they are not intended to prove discrimination, are useful for tracking potentially discriminatory decision-making. Given the counterfactual model of discrimination and our focus on the comparator, the main fairness works useful for discrimination are counterfactual fairness (Kusner et al., 2017) and individual fairness (Dwork et al., 2012). The former establishes the same treatment of the factual and its counterfactual instance under the same decision; the latter establishes the same treatment of similar instances under the same decision. Under the counterfactual model of discrimination, these two concepts intersect as we use the similarity between individuals to approximate the comparison between an individual and its counterfactual-self. Works like the FlipTest (Black et al., 2020) or counterfactual situation testing (Álvarez and Ruggieri, 2023) build upon these two definitions. Overall, this holds for other similar works including (Dutta et al., 2021; Anthis and Veitch, 2023; Salimi et al., 2019).
We view algorithmic fairness tools based on generative models as central to implementing the MM comparator. It includes that are not necessarily concerned with discrimination testing but on, e.g., learning fair representations (Zemel et al., 2013). Tools that are more focused on proving discrimination, meaning along the same lines as discrimination discovery, still face the key challenge of linking unfairness to discrimination (Weerts et al., 2023). These tools are useful, but it remains to be seen how they fit into the pipeline for establishing discrimination. This is not a technical challenge in-itself, but one on whether the technology is embraced or not by the relevant stakeholders. Another key challenge, which links directly to the MM comparator, is how these tools address indirect discrimination as most algorithmic decision-makers are not allowed to use the protected attribute as input. As illustrated in Section 4, the choice of comparator affects the number of discrimination cases identified. These tools, along with the discrimination discovery ones, are capable of generating more complex comparators than the standard tools. The question then is whether they should aim or not at implementing the MM comparator over the CP comparator. Our view is, at a minimum, they should allow for the possibility of choosing between the two.
6. Conclusion
We discussed extensively the problem of discrimination testing by focusing on the notion of comparator. We presented the derivation of the comparator as a causal modeling problem, emphasizing the role ML methods like generative models can play for better discrimination testing methods. We introduced the ceteris paribus (CP) comparator and mutatis mutandis (MM) comparator distinction; the former refers to the standard comparator and the latter represents a new kind of comparator that departs from idealized comparisons. Experiments showcase the impact this distinction has on discrimination testing and motivates the need to consider, at least, the option of being able to choose the kind of comparator when implementing a tool for discrimination testing.
This work comes from a ML perspective on discrimination testing. We have done our best in incorporating other fields and their views on the need for the MM comparator. Future work should revisit the problem of discrimination testing by considering the full legal procedure for establishing discrimination. Future work should consider other generative models for implementing the MM comparator beyond SCM. We are certain that the MM comparator and its departure from the CP comparator, both as a conceptual and modeling problem, will motivate further fair ML work on discrimination testing.
Acknowledgements.
This work has received funding from the European Union’s Horizon 2020 research and innovation program for the MSCA project NoBIAS (g.a. No. 860630) and for the project FINDHR (g.a. No. 101070212). This work reflects only the authors’ views and the European Research Executive Agency is not responsible for any use that may be made of the information it contains.

### v3

@@
 Abstract.
-Testing for discrimination consists of deriving a profile, known as the comparator, similar to the profile making the discrimination claim, known as the complainant, and comparing the outcomes of these two profiles.
-An important aspect for establishing discrimination is evidence, often obtained via discrimination testing tools that implement the complainant-comparator pair.
-In this work, we revisit the role of the comparator in discrimination testing.
-We argue for the causal modeling nature of deriving the comparator, and introduce a two-kinds classification for the comparator: the ceteris paribus (CP), and mutatis mutandis (MM) comparators.
-The CP comparator is the standard one among discrimination testing, representing an idealized comparison as it aims for having a complainant-comparator pair that only differs on membership to the protected attribute.
-As an alternative to it, we define the MM comparator, which requires that the comparator represents what would have been of the complainant without the effects of the protected attribute on the non-protected attributes.
-The complainant-comparator pair, in that case, may also be dissimilar in terms of all attributes.
-We illustrate these two comparators and their impact on discrimination testing using a real illustrative example.
-Importantly, we position generative models and, overall, machine learning methods as useful tools for constructing the MM comparator and, in turn, achieving more complex and realistic comparisons when testing for discrimination.
-algorithmic fairness, discrimination discovery, counterfactual model of discrimination
+Testing for individual discrimination involves deriving a profile, the comparator, similar to the one making the discrimination claim, the complainant, based on a protected attribute, such as race or gender, and comparing their decision outcomes.
+The complainant-comparator pair is central to discrimination testing.
+Most discrimination testing tools rely on this pair to establish evidence for discrimination.
+In this work we revisit the role of the comparator in discrimination testing.
+We first argue for the inherent causal modeling nature of deriving the comparator.
+We then introduce a two-kinds classification for the comparator: the ceteris paribus, or “with all else equal,” (CP) comparator and the mutatis mutandis, or “with the appropriate adjustments being made,” (MM) comparator.
+The CP comparator is the standard comparator, representing an idealized comparison for establishing discrimination as it aims for a complainant-comparator pair that only differs on membership to the protected attribute.
+As an alternative to it, we define the MM comparator, which requires that the comparator represents the “what would have been of” the complainant without the effects of the protected attribute on the non-protected attributes.
+Under the MM comparator, the complainant-comparator pair can be dissimilar in terms of the non-protected attributes, departing from an idealized comparison.
+Notably, the MM comparator is a more complex kind of comparator and its implementation offers an impactful venue for machine learning methods.
+We illustrate these two comparators and their impact on discrimination testing using a real-world example.
+algorithmic fairness, discrimination discovery, counterfactual reasoning
 1. Introduction
-The concept of discrimination refers to an unjustified difference in treatment toward an individual or group of individuals based on (perceived) membership to a protected-by-law group (Romei and Ruggieri, 2014).
-We view, thus, discrimination as unfairness sanctioned by law.
-Testing for discrimination against an individual instance is a comparative process centered around the complainant-comparator pair.
-As the names would suggest, the complainant is the individual instance motivating the discrimination claim in court, while the comparator is the individual instance used for validating the claim.
-We use the comparator to answer the hypothetical question behind any discrimination claim: would the decision have been the same if the complainant had been a member of the non-protected group instead of the protected group under non-discrimination law?
+Discrimination refers to an unjustified difference in treatment toward an individual based on (perceived) membership to a protected by non-discrimination law group (Romei and Ruggieri, 2014).
+Testing for discrimination is a comparative process centered around the complainant-comparator pair.
+As the names would suggest, the complainant is the individual instance motivating the discrimination claim while the comparator is the individual instance used for validating it.
+We use the comparator to answer the hypothetical question behind all discrimination claims: would the decision have been the same if the complainant had been a member of the non-protected rather than the protected group?
 The comparison between the outcomes of the complainant and its comparator helps establish whether the decision-making process in question is or not discriminatory.
 Most tools for discrimination testing, from the traditional methods like correspondence studies (Bertrand and Mullainathan, 2004) and situation testing (Bendick, 2007) to the recent algorithmic methods like discrimination discovery (Ruggieri et al., 2010) and counterfactual fairness (Kusner et al., 2017), rely on this pair.
-Importantly, these tools not only implement the complainant-comparator comparison but also derive the complainant’s comparator as a whole.
+These tools not only implement the complainant-comparator comparison but also derive the complainant’s comparator as a whole based on the relevant process data.
+The comparator is central to establishing evidence for discrimination.
 Non-discrimination law requires similar, or similarly situated, individuals to be treated similarly regardless of gender, race, or other protected attributes.
-It is a requirement largely shared across countries, only with some variation on the list of protected attributes (Romei and Ruggieri, 2014).
-The comparator is meant to represent a similar (or, at least, comparable) individual to the complainant, allowing to envision an alternative path for the complainant conditional on “changing” its protected attribute.
-It requires what we view, in practice, as an imaginative act in which we find (e.g., via matching models (Qureshi et al., 2020)) or create (e.g., via generative models (Black et al., 2020)) an individual instance representative enough of the complainant’s alternative world.
-In fact, the comparator, unlike the complainant, can also be a fictitious individual instance (Weerts et al., 2023).
-In a time of modern decision flows in which machine learning (ML) methods are increasingly used to aid as well as regulate decision-making processes, such modeling flexibility around the comparator offers a promising ML application.
-The choice of comparator, meaning how these tools for discrimination testing derive the comparator, matters for three incremental reasons.
-First, it embodies a normative statement on what similarity between individuals is and, in turn, how the protected attribute influences individual outcomes within the context of interest (e.g., (Weerts et al., 2023)).
-Second, it conditions the results of these tools by determining which individual outcomes are to be compared for evaluating the discrimination claim (e.g., (Álvarez and Ruggieri, 2023)).
-Third, it acts as a target for the sort of substantive equality between individuals that we wish to achieve as a society through non-discrimination law (e.g., (Wachter et al., 2020)).
+It is a requirement shared across countries with only some variation on the list of protected attributes (Romei and Ruggieri, 2014).
+The comparator is meant to represent a similar individual to the complainant and allows to envision an alternative path for the complainant conditional on “changing” its protected attribute.
+It requires what we view in practice as an imaginative act in which we find (using, for instance, matching models (Qureshi et al., 2020)) or create (using generative models (Black et al., 2020)) an individual instance representative enough of the complainant’s alternative world.
+In fact, the comparator, unlike the complainant that motivates it, can be a fictitious individual instance (Weerts et al., 2023).
+The choice of comparator, meaning how these discrimination testing tools derive the comparator, matters for three incremental reasons.
+First, it embodies a normative statement on what similarity between individuals is and, in turn, how the protected attribute influences individual outcomes within the context of interest (Weerts et al., 2023).
+Second, it conditions the results of these tools by determining which individual outcomes are to be compared for evaluating the discrimination claim (Álvarez and Ruggieri, 2025).
+Third, it acts as a target for the sort of substantive equality between individuals that we wish to achieve as a society through non-discrimination law (Wachter et al., 2020).
 Most tools aim to define the comparator as an observably similar instance to the complainant all but for membership to the protected attribute, such as defining a male candidate with the same work experience as the female complainant in a hiring decision or a white applicant with the same credit score as the non-white complainant in a lending decision.
 We refer to this standard operationalization of the comparator as the ceteris paribus, or “with all else being equal,” (CP) comparator.
-Precisely because of these reasons, debates around the CP comparator predate modern decision flows, e.g., (Westen, 1982; Heckman, 1998; Kohler-Hausmann, 2019), and remain relevant today as fair ML researchers tackle the new problem of algorithmic discrimination, e.g., (Kleinberg et al., 2019; Ruggieri et al., 2023; Álvarez et al., 2024).
-We would add that this reliance on the CP comparator is problematic for fair ML.
-It overlooks the potential positive role ML methods, particularly generative models, could play as tools for deriving more complex comparators that address the ongoing debates.
+Precisely because of these three reasons, debates around the CP comparator predate modern decision flows that rely on machine learning (ML) (Westen, 1982; Heckman, 1998; Kohler-Hausmann, 2019).
+These debates remain relevant as ML researchers tackle the new problem of algorithmic discrimination (Kleinberg et al., 2019; Ruggieri et al., 2023; Álvarez et al., 2024).
 This work studies the problem of testing for discrimination by revisiting the role of the comparator.
-We do so by, first, embracing the causal modeling nature of discrimination testing and, second, proposing an alternative to the CP comparator.
+We do so by, first, embracing the causal modeling nature of discrimination testing and, second, proposing an alternative to the standard CP comparator.
 Regarding the first point, given that the comparator answers to the “what would have been if” question around the complainant, we assert that the comparator is a counterfactual representation of the factual complainant.
-We further assert that the derivation of the comparator itself constitutes a causal modeling problem and that such problem formulation is, for now, the best possible one given the comparative nature of discrimination testing.
+We further assert that the derivation of the comparator itself constitutes a causal modeling problem and that such problem formulation is the best possible one given the comparative nature of discrimination testing.
 Regarding the second point, we propose the mutatis mutandis, or “with the appropriate adjustments being made,” (MM) comparator.
 The MM comparator requires that all non-protected attributes constituting the comparator account for the downstream effects of the protected attribute.
-The CP comparator, thus, represents an idealized comparison to the complainant, while the MM comparator represents an adjusted comparison to the complainant.
-With the MM comparator, we expand the problem of discrimination testing and, given the greater complexity of deriving a MM comparator relative to a CP comparator (read, a stronger “imagination”), we identify an impactful application for fair ML models.
-Example 1.1 below motivates and illustrates the difference between the two comparators.
+Hence, the CP comparator represents an idealized comparison to the complainant while the MM comparator represents an adjusted comparison to the complainant with each based on how we conceive modeling-wise the influence of the protected attribute on all other attributes.
+Example 1.1 motivates and illustrates the difference between the two comparators.
+Altogether, we view the implementation of the MM comparator as an impactful application for ML models.
+Implementing the MM comparator relative to the CP comparator imposes additional complexity (read, a stronger imaginative act) to the problem of discrimination testing that can be better addressed by ML models over more traditional models.
+This additional point contrasts with popular works, mainly within the fair ML field (Kohler-Hausmann, 2019; Hu and Kohler-Hausmann, 2020), that are adamant toward treating discrimination testing as a ML problem.
 Example 1.1.
-(Tenure at U) This example is based on Morgan et al.
+(Tenure at U)  [^1] This example is based on Morgan et al.
 (2021)’s research on the unequal impact of parenthood in academia for female researchers.
-[^1] Social expectations and household dynamics are key factors in creating this difference.
-Mothers, as the primary caregivers, tend to carry the burden of childbearing within their households, which consumes potential time that could be spent on advancing ones academic career by, e.g., writing papers at home or networking at conferences.
+Social expectations and household dynamics are key factors in creating this difference.
+Mothers, as the primary caregivers, tend to carry the burden of childbearing within their households, which consumes potential time that could be spent on advancing ones academic career by, for instance, writing papers at home or networking at conferences.
 Fathers tend to share less of this burden.
@@
 (2021) for details.
-Also see Roman and Rodríguez (2023).
+Also see Roman and Rodríguez (2023) for similar results in couples outside of academia.
 Let us assume that the university U only considers the number of publications for granting tenure.
@@
 Assuming some quantifiable penalty on the number of publications from motherhood, would we be willing to use Vincent, with 18 publications and recently tenured, as Clara’s comparator instead of Mike?
-In Example 1.1 Mike represents the CP comparator (Definition 2.1) while Vincent the MM comparator (Definition 3.1).
+In Example 1.1, Mike represents the CP comparator (Definition 2.1) while Vincent represents the MM comparator (Definition 3.1).
 Depending on which one we use, assuming a literal comparison, we would invalidate or validate, respectively, Clara’s discrimination claim.
@@
 Given what we know from works like Morgan et al.
-(2021), we are implicitly asking which male researcher is more similar to Clara?
-It is not an easy question to answer, highlighting the complexity behind modeling the comparator.
+(2021), in this example we are implicitly asking which male researcher is more similar to Clara given what we know about gender and its role in this context?
+We believe that it is not an easy question to answer, highlighting the complexity behind modeling the comparator.
 With Example 1.1 and, overall, with this work we are not suggesting that the CP comparator is wrong or that the MM comparator is better.
-Instead, we want to show that the MM comparator is an alternative formulation to the CP comparator and, overall, stress the importance of the choice of comparator.
-Policies like non-discrimination law and positive actions aimed at addressing/removing systematic inequalities require concrete targets (Wachter et al., 2020).
-Discrimination testing tools, essentially, enforce such targets through the comparator, with the CP and MM comparators representing two distinct targets.
+We simply want to position the MM comparator as a viable alternative formulation to the standard CP comparator and, in doing so, stress the importance of the choice of comparator in discrimination testing.
+Policies like non-discrimination law and positive actions aimed at addressing systematic inequalities require concrete targets (Wachter et al., 2020).
+Discrimination testing tools enforce such targets through the comparator.
+The CP and MM comparators, thus, represent two distinct targets.
+It is a missed opportunity to rely solely on the CP comparator, especially, when we already have ML modeling tools able to implement the more complex MM comparator.
+In this work, we focus on one of such ML tools by using structural causal models (SCM) Pearl (2009) to derive the MM comparator.
+Contributions.
+Our main contributions are twofold.
+First, we demonstrate why the derivation of the comparator is best understood as a causal modeling problem centered on counterfactual reasoning.
+This is indeed a divisive view on establishing discrimination but one that should be embraced for developing better discrimination testing tools.
+Second, with the MM comparator we propose a new approach to the derivation of the comparator and introduce the distinction between CP and MM comparators.
+This novel distinction expands the methodological landscape for establishing discrimination.
+Additionally, our illustrative experiment showcases how this distinction in comparators impacts discrimination testing in practice.
+The need for implementing the more imaginative MM comparator clearly positions ML models as promising approaches within discrimination testing.
 Structure.
 In the remainder of this section we discuss our modeling choice for causality (Section 1.1) and relevant related work (Section 1.2).
-Moving forward, in Section 2 we argue that deriving the comparator is a causal modeling problem and formulate the CP comparator, which is the standard in discrimination testing.
+In Section 2 we argue that deriving the comparator is a causal modeling problem and formulate the CP comparator.
 In Section 3, we propose the MM comparator and justify its departure from the CP comparator.
-In Section 4, using a popular real dataset in fair ML, we showcase both comparators by comparing two related discrimination testing tools—situation testing (Thanh et al., 2011) and counterfactual situation testing (Álvarez and Ruggieri, 2023)—that implement, respectively, the CP comparator and MM comparator.
+In Section 4, using a popular real dataset in fair ML, we showcase both comparators by comparing two related discrimination testing tools—situation testing (Thanh et al., 2011) and counterfactual situation testing (Álvarez and Ruggieri, 2023, 2025)—that implement, respectively, the CP comparator and MM comparator.
 In Section 5, we provide a representative survey of discrimination testing tools.
-In Section 6, we conclude.
-Contributions.
-Our main contributions are threefold.
-First, we demonstrate why the derivation of the comparator is best understood as a causal modeling problem.
-Second, with the MM comparator we propose a new approach to deriving the comparator and, in turn, introduce the distinction between CP and MM comparators.
-Our experiments illustrate how this distinction impacts discrimination testing.
-Third, we position ML methods as a promising tool for achieving the MM comparator and developing better tools for discrimination testing.
-Audience and scope.
+In Section 6, we conclude this work.
+Audience, scope, and word choice.
 Discrimination testing is an interdisciplinary topic.
-Our work is aimed primarily at ML researchers working on this topic and, hence, our framing of it as a modeling problem.
-We are interested in quantitative tools used for deriving prima facie evidence for discrimination claims, which constitutes one aspect of the legal process for uncovering discrimination.
+Our work is aimed primarily at ML researchers.
+We are interested in the quantitative tools used for deriving prima facie evidence for discrimination claims, which constitutes one of many aspects of the legal process for establishing discrimination.
+[^2] Technically, all discrimination tools can only test for prima facie discrimination (Romei and Ruggieri, 2014; Weerts et al., 2023; Álvarez et al., 2024).
+This distinction is sometimes downplayed in the corresponding papers.
+We come back to this distinction in Section 2.2.
+Throughout the paper we focus on individual discrimination.
+Namely, we are interested at testing whether an individual instance of a protected group has been discriminated.
+The notion of group discrimination tackles instead discrimination against a subset of individuals from a protected group.
+[^3] However, Binns (2020) argues that the apparent conflict between individual and group fairness is an artifact of fairness metrics rather than reflecting different normative principles.
+Here, we stick with the current fair ML convention.
+Notably, although often used interchangeably, we make a clear distinction between unfair and discriminatory decisions.
+Both concepts focus on judging the decision outcomes around a specific attribute; however, discrimination has a stronger meaning than unfairness as it requires such attribute to be under non-discrimination law.
+Unfairness, in that sense, is more flexible in meaning.
+All discriminatory decisions are unfair, but not all unfair decisions are discriminatory.
+Discrimination, in short, can be viewed as unfairness sanctioned by law.
 Our focus is mainly on the operationalization of the comparator, which we view as a ML problem that holds for any modern decision flow, regardless of the nature of the decision-maker.
 Throughout the paper, we use generative model to refer to a model capable of creating synthetic data (McElreath, 2018).
-Similarly, we use counterfactuals to refer to interventional probability distributions obtained through causal reasoning (Pearl, 2009), not to be confused with counterfactual explanations222 Counterfactual explanations refer to individual instances similar to the one that is explained but have a different outcome.
+These are models capable of learning enough about the data generating process of interest to generate instances that are both novel and conceivable under said process, which includes, by definition, approaches like representation learning (Zemel et al., 2013) and SCM Pearl (2009).
+Similarly, we use counterfactuals to refer to interventional probability distributions obtained through causal reasoning (Pearl, 2009), not to be confused with counterfactual explanations [^4] Counterfactual explanations refer to individual instances similar to the one that is explained but have a different outcome.
 To avoid confusion, Miller (2019) proposes to name them contrastive explanations, and reserve the term counterfactual as understood in the statistical causality literature.
@@
 1.1. Auxiliary Causal Knowledge
-We formulate causality using structural causal models (SCM) (Pearl, 2009), which is the most used framework within ML.
+We formulate causality using structural causal models (SCM) (Pearl, 2009), a popular framework within causal ML.
 SCM are probabilistic graphical models that link (latent) exogenous variables to (observable) endogenous variables via structural equations.
@@
 Formally, a SCM $\mathcal{M}$ is a tuple $\mathcal{M}=\langle\mathbf{U},\mathbf{X},\mathbf{F}\rangle$, where $\mathbf{U}=(U_{1},\ldots,U_{p})$ are $p$ independent exogenous latent variables, with $\mathbf{U}\sim P(\mathbf{U})=\prod_{j}P(U_{j})$, $\mathbf{X}=(X_{1},\ldots,X_{p})$ are $p$ endogenous observed variables, and $\mathbf{F}=(f_{1},\ldots,f_{p})$ are structural equations such that:
-(1) $\text{for $j=1,\dots,p$}\;\;\;X_{j}:=f_{j}(X_{pa(j)},U_{j})$
-Each $f_{j}$ defines the observed variable $X_{j}$ on the basis of its causal parents $X_{pa(j)}$ and of the latent variable $U_{j}$ Associated to the SCM $\mathcal{M}$ is the causal graph $\mathcal{G}$, in which each node is a random variable and the directed edges between them a causal effect.
+(1) $\text{for $j=1,\dots,p$}\;\;\;X_{j}:=f_{j}(X_{pa(j)},U_{j}).$
+Each $f_{j}$ defines the observed variable $X_{j}$ on the basis of its causal parents $X_{pa(j)}$ and of the latent variable $U_{j}$ .
+Associated to the SCM $\mathcal{M}$ is the causal graph $\mathcal{G}$, in which each node is a random variable and the directed edges between them a causal effect.
 We assume a directed acyclical graph, or DAG, $\mathcal{G}$.
 Through $\mathcal{G}$ we represent how the variables relate to each other in terms of cause-effect pairs.
-For instance, the causal graph $X_{1}\rightarrow X_{2}$ reads as “$X_{1}$ causes $X_{2}$”.
-The SCM $\mathcal{M}$ in (1) can act as a generative model.
-In particular, it can generate counterfactual distributions that answer to counterfactual questions as the ones asked in discrimination claims.
+The causal graph $X_{1}\rightarrow X_{2}$, for instance, reads as $X_{1}$ causes $X_{2}$”.
+The SCM $\mathcal{M}$ in (1) acts as a generative model by generating counterfactual distributions that answer to counterfactual questions as the ones asked in discrimination claims.
 Let $A$ be one of the variables in $\mathbf{X}$ denoting the protected attribute.
@@
 Prediction: we generate the counterfactual distribution $P(\mathbf{X}^{CF}_{A\leftarrow\alpha}\;|\;\mathbf{X})$ by propagating the abducted $P(U_{j}\;|\;\mathbf{X})$ through the revised structural equations in $\mathcal{M}^{\prime}$.
-Beyond Section 4, in which we rely on counterfactual generation to implement the MM comparator, what we discuss in Sections 2 and 3 is not exclusive to SCM.
-It is our generative model of choice, mainly because previous works on discrimination testing (e.g., (Kusner et al., 2017; Kilbertus et al., 2017; Álvarez and Ruggieri, 2023)) use it.
-Here, our goal is not prioritize one generative model over the others, but to focus on one generative model to show how the others could be useful for the problem of discrimination testing.
-1.2. Relevant Related Work
+Importantly, beyond Section 4, in which we rely on counterfactual generation to implement the MM comparator, what we discuss in Sections 2 and 3 is not exclusive to SCM.
+It is our generative model of choice, mainly because previous works on discrimination testing (Kusner et al., 2017; Kilbertus et al., 2017; Álvarez and Ruggieri, 2023, 2025) use it.
+Our goal is not prioritize one generative modeling approach over the others, but to focus on one generative model approach to show how the others could be useful for the problem of discrimination testing under the MM comparator.
+1.2. Related Work
 The literature on discrimination testing is vast, multidisciplinary, and regional.
-We refer to (Chen et al., 2024) for a recent survey on testing fairness of ML models, and to (Hutchinson and Mitchell, 2019; Romei and Ruggieri, 2014) for surveys with a multidisciplinary perspective.
-For a general overview of bias and fairness in AI, see (Álvarez et al., 2024; Ntoutsi et al., 2020; Mehrabi et al., 2022).
-Throughout the paper, we consider individual discrimination.
-Namely, we are interested at testing whether an individual instance of a protected group has been discriminated.
-The notion of group discrimination tackles instead discrimination against a subset of individuals from a protected group.
-However, Binns (2020) argues that the apparent conflict between individual and group fairness is an artifact of fairness metrics rather than reflecting different normative principles.
-In this subsection, we position our work relative to three recent papers that we view as directly applicable to the modeling problem of revisiting the comparator.
+We refer to Chen et al.
+(2024) for a recent survey on testing fairness of ML models, and to Hutchinson and Mitchell (2019) and Romei and Ruggieri (2014) for surveys with a multidisciplinary perspective.
+For a general overview of bias and fairness in AI, see Álvarez et al.
+(2024); Ntoutsi et al.
+(2020) and Mehrabi et al.
+(2022).
+Here, we wish to position our work with respect to three papers that are directly applicable to the modeling problem of revisiting the comparator in discrimination testing.
 To the best of our knowledge, we are the first to explicitly tackle this problem.
-Later in Section 5, we provide a representative survey of discrimination testing tools, with a focus on the comparator chosen by those tools.
-We do so as we first want to introduce the CP and MM comparators (Sections 2 and 3), and illustrate their impact on a discrimination testing problem (Section 4).
+Later in Section 5, we provide a selective survey of discrimination testing tools, with a focus on the comparator chosen by those tools.
+We do so as we first want to introduce the CP and MM comparators in Sections 2 and 3, and illustrate their impact on a discrimination testing problem in Section 4.
 First, Kohler-Hausmann (2019) argues that discrimination testing focuses on answering counterfactual questions about the complainant.
-She refers to the current approach as the causal counterfactual model of discrimination, and states that it is flawed due to its treatment of protected attributes.
-In particular, using a constructivist approach to race, she argues that comparing similar individuals that only differ on skin color fails to capture how race affects all other aspects of these individuals’ lives.
-We draw considerably from Kohler-Hausmann (2019) as she writes the first critique to the CP comparator.
+The work refers to the current approach as the causal counterfactual model of discrimination, and states that it is flawed due to its treatment of protected attributes.
+In particular, using a constructivist approach to race, the work argues that comparing similar individuals that only differ on skin color fails to capture how race affects all other aspects of these individuals’ lives.
+We draw considerably from Kohler-Hausmann (2019) as it represents the first critique to the CP comparator.
 However, different from Kohler-Hausmann (2019), we propose an alternative to the CP comparator in the form of the MM comparator and embrace the role of counterfactual reasoning within discrimination testing.
-This last point is important as Kohler-Hausmann’s view (see, e.g., Hu and Kohler-Hausmann (2020)) is largely against causal modeling for discrimination testing and is suspicious toward attempts to measure the so-called causal effect of race.
-She fails, though, to provide an alternative to the current model.
-We disagree with her position toward causal modeling.
-What we agree with her is that we can (and, potentially, should) do better than the CP comparator.
-The MM comparator is a response to Kohler-Hausmann (2019).
+This last point is important as Kohler-Hausmann’s view (see, for instance, Hu and Kohler-Hausmann (2020)) is largely against causal modeling for discrimination testing and is suspicious toward attempts to measure the so-called causal effect of race.
+Kohler-Hausmann fails, though, to provide an alternative to the current model.
+We disagree with Kohler-Hausmann (2019)’s position toward causal modeling, and agree that we can and, potentially, should do better than the CP comparator.
+The MM comparator is our response to Kohler-Hausmann (2019).
 Second, Wachter et al.
-(2020) state that non-discrimination law aims for substantive equality, meaning it should aim at changing the status quo of the societal context of interest.
-[^3] Their argument is specific to the European Union.
+(2020) states that non-discrimination law aims for substantive equality, meaning it should aim at changing the status quo of the societal context of interest.
+[^5] The argument is specific to the European Union.
 Although it seems not to be the case for US non-discrimination law (Barocas and Selbst, 2016), the goal of substantive equality is implicit in many of the fair ML definitions used by US researchers as argued by Wachter et al.
 (2020).
-Under this premise, they classify fair ML methods into bias preserving and bias transforming based on whether they assume (the former) or not (the latter) a neutral status quo.
+Under this premise, the work classifies fair ML methods into bias preserving and bias transforming based on whether they assume (the former) or not (the latter) a neutral status quo.
 Although Wachter et al.
-(2020) make a strong case for ML methods as societal shapers, they lack clarity on how these methods are to be implemented, specially, with respect to non-discrimination law.
+(2020) makes a strong case for ML methods as societal shapers, the work lacks clarity on how these methods are to be implemented, specially, with respect to non-discrimination law.
 We tackle this gap by formalizing the MM comparator.
@@
 (2020) is that non-discrimination law must not aim at reaching equality of outcomes only, but also at imposing equality of opportunities in a systematic way.
-Such a step requires envisioning an ideal target, which needs to be generated, for the current context to aim at.
+Such a step requires envisioning an ideal target, which needs to be generated for the current context to aim at.
 The MM comparator represents the goal of substantive equality in the context of discrimination testing.
-Third, Álvarez and Ruggieri (2023) introduce a similar classification to the CP and MM comparators distinction.
-Drawing from Kohler-Hausmann (2019)’s discussion on the causal counterfactual model of discrimination, they state that discrimination testing tools can be broadly classified into comparators that do and do not consider the downstream effects of the protected attribute on all other attributes.
-They offer a valuable first step to the goal of revisiting the comparator, but fall short in formalizing their claims on the comparator.
-They neither define these kinds of comparators nor they survey the quantitative tools to illustrate them.
-On top of addressing these shortcoming from Álvarez and Ruggieri (2023), we discuss the role of counterfactual reasoning in discrimination testing, which they take for granted.
+Third, Álvarez and Ruggieri (2023) introduces the CP and MM classification but briefly and without the necessary discussion.
+[^6] The same criticism applies to the journal version: Álvarez and Ruggieri (2025).
+We do this on purpose.
+The CP versus MM discussion merits its own work as it applies to all discrimination testing tools.
+This is because the focus of the paper is on formulating the discrimination testing method of counterfactual situation testing, which is a possible implementation of the MM comparator.
+Drawing from Kohler-Hausmann (2019)’s discussion on the causal counterfactual model of discrimination, the work states that discrimination testing tools can be broadly classified into comparators that do and do not consider the downstream effects of the protected attribute on all other attributes.
+The work offers a valuable first step to the goal of revisiting the comparator, but falls short in formalizing the claims on the comparator.
+The work neither defines these kinds of comparators nor surveys the quantitative tools to illustrate them.
+On top of addressing these shortcoming from Álvarez and Ruggieri (2023), we discuss the role of counterfactual reasoning in discrimination testing, which the works takes for granted.
 We note that the causal treatment of discrimination, specifically, using SCM is a common practice.
-See (Makhlouf et al., 2024) for a survey.
-In particular, Kilbertus et al.
+See Makhlouf et al.
+(2024) for a survey.
+Kilbertus et al.
 (2017) are the first to formulate problems like direct, indirect, and proxy discrimination using SCM, while Plecko and Bareinboim (2024) try to summarize discrimination within a general causal fairness model using SCM.
 Works like these are relevant, but lack interdisciplinary depth.
-For instance, both papers fail to discuss non-discrimination law and how exactly the modeling tools based on their formulations can be (or need to be) used as evidence.
+For instance, both papers fail to discuss non-discrimination law and how exactly the modeling tools based on their formulations can be or need to be used as evidence.
 Their treatment of the subject is detached from the real-world application.
-These papers, e.g., do not discuss the comparator.
+These papers do not discuss the comparator.
 To be clear, such papers have a different scope—mainly to formulate discrimination using SCM—but, because of such approach, fail to address the role of causality in discrimination testing.
-[^4] Such a discussion is more explicit in the social sciences, in which causal models are also used for testing discrimination.
-See, e.g., Heckman (1998).
-In that sense, our work complements these and similar works by revisiting the comparator.
+Such a discussion is more explicit in the social sciences, in which causal models are also used for testing discrimination.
+See, for instance, Heckman (1998).
+Our work complements these and similar works by revisiting the comparator.
 2. The Comparator as a Causal Modeling Problem
-Testing for discrimination amounts to a two-step modeling problem centered on deriving the comparator $i^{\prime}$, and comparing it to the discrimination claim’s complainant $i$.
-In this section, we focus on the derivation of the comparator, motivating the need for $i^{\prime}$ and its inherent challenges.
-Importantly, we argue that the derivation of the comparator is a causal modeling problem not only because we use the comparator to demonstrate whether the decision is or not caused by the protected attribute, but also because $i^{\prime}$ is a counterfactual representation of the factual $i$.
+In this section, we argue that testing for discrimination amounts to a two-step modeling problem centered on deriving the comparator $i^{\prime}$ and comparing it to the discrimination claim’s complainant $i$.
+We argue for the derivation of the comparator, motivating the need for $i^{\prime}$.
+We also argue that the derivation of the comparator is a causal modeling problem not only because we use $i^{\prime}$ to demonstrate whether the decision is or not caused by the protected attribute, but also because $i^{\prime}$ is a counterfactual representation of the factual $i$.
 2.1. The Setup
-Non-discrimination law mandates that a decision-maker—be it algorithmic or human—treats similar individuals similarly regardless of the protected attribute (Kleinberg et al., 2019).
+Non-discrimination law mandates that a decision-maker, be it algorithmic or human, treats similar individuals similarly regardless of the protected attribute (Kleinberg et al., 2019).
 Discrimination testing tools uncover discrimination by finding and comparing similar individuals to isolate the effect of the protected attribute on the decision outcome (Romei and Ruggieri, 2014).
@@
 Definition 2.1.
-(The CP Comparator) Let the individual profile $\langle\mathbf{x}_{i},a_{i},y_{i}\rangle$ represent the complainant $i$ on which the discrimination claim is based.
-We define a ceteris paribus (CP) comparator $i^{\prime}$ as an individual profile $\langle\mathbf{x}_{i^{\prime}},a_{i^{\prime}},y_{i^{\prime}}\rangle$ such that:
+(The CP Comparator) Let the real individual profile $\langle\mathbf{x}_{i},a_{i},y_{i}\rangle$ represent the complainant $i$ on which the discrimination claim is based on.
+We define the ceteris paribus (CP) comparator $i^{\prime}$ as a real or fictitious individual profile $\langle\mathbf{x}_{i^{\prime}},a_{i^{\prime}},y_{i^{\prime}}\rangle$ such that:
 (2) $d\big(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\big)\leq\epsilon\;\;\;\text{and}\;\;\;a_{i}\neq a_{i^{\prime}}$
-where $d$ is a distance function [^5] The issue of engineering fairness is, without doubts, challenging (Scantamburlo, 2021).
-In this regard, we believe that the definition of the distance function $d$ is not only a technical choice (see the survey (Blanco-Mallo et al., 2023)), but is necessarily specific of the application domain at hand.
-For instance, in job screening, $d$ can result from a weighting of the skills of job candidates: how to set the weights is a non-trivial issue, which requires an agreement among multiple stakeholders.
-For this reason, in this paper, we make no assumption/preference about the definition of $d$.
-(the closer to 00 the more similar are $\mathbf{x}_{i}$ and $\mathbf{x}_{i^{\prime}}$) and ϵ+italic-$\epsilon\in\mathbb{R}^{+}$ a threshold.
-Importantly, it is possible to provide a hypothetical comparator (Weerts et al., 2023), meaning we may find or generate an $i^{\prime}$ similar under $d$ to $i$.
-The comparator derived from (2) is the standard comparator in discrimination testing (Heckman, 1998; Kohler-Hausmann, 2019).
-[^6] There is a conception of discrimination testing that is non-comparative, though it is not widely used.
-See Binns (2020).
+where $d(\cdot,\cdot)$ is a distance function [^7] The issue of engineering fairness is, without doubts, challenging (Scantamburlo, 2021).
+We believe that the definition of the distance function $d(\cdot,\cdot)$ is not only a technical choice (see the survey by Blanco-Mallo et al.
+(2023)), but is necessarily specific of the application domain at hand.
+For instance, in job screening, $d(\cdot,\cdot)$ can result from a weighting of the skills of job candidates: how to set the weights is a non-trivial issue, which requires an agreement among multiple stakeholders.
+For this reason, in this paper, we make no assumption/preference about the definition of $d(\cdot,\cdot)$.
+and $\epsilon\in\mathbb{R}^{+}$ a threshold.
+The closer $d(\cdot,\cdot)$ is to 0 the more similar are $\mathbf{x}_{i}$ and $\mathbf{x}_{i^{\prime}}$.
+We emphasize that it is possible to provide a hypothetical comparator (Weerts et al., 2023), meaning we may find or generate an $i^{\prime}$ similar under $d$ to $i$.
+The comparator derived from Equation (2) is the standard comparator in discrimination testing (Heckman, 1998; Kohler-Hausmann, 2019).
+[^8] There is a conception of discrimination testing that is non-comparative, though it is not widely used (Binns, 2020).
 We refer to it as the ceteris paribus, or “with all else being equal,” (CP) comparator as, ideally, we expect to find an $i^{\prime}$ profile all equal but for the protected attribute to the $i$ profile when $\epsilon=0$.
@@
 Definition 2.2.
-(Difference in Outcomes) Given the individual profiles $\langle\mathbf{x}_{i},a_{i},y_{i}\rangle$ and $\langle\mathbf{x}_{i^{\prime}},a_{i^{\prime}},y_{i^{\prime}}\rangle$, we test the discrimination claim by focusing on the difference in outcomes:
+(Difference in Outcomes) Given the individual profiles $\langle\mathbf{x}_{i},a_{i},y_{i}\rangle$ and $\langle\mathbf{x}_{i^{\prime}},a_{i^{\prime}},y_{i^{\prime}}\rangle$, we test the discrimination claim by focusing on the difference in a same decision outcome:
 (3) $\delta=P(Y=y_{i^{\prime}}\,|\,\mathbf{X}=\mathbf{x}_{i^{\prime}},A=a_{i^{\prime}})-P(Y=y_{i}\,|\,\mathbf{X}=\mathbf{x}_{i},A=a_{i})$
-where, $\delta$ may, e.g., be greater or equal to 0 or need to follow a known distribution depending on the legal context.
+where, $\delta$ may, for example, be greater or equal to 0 or need to follow a known distribution depending on the legal context.
+Notably, we focus on the same decision between the complainant and the comparator, meaning $y_{i^{\prime}}=y_{i}$.
 The difference in outcomes (3) is not based on a literal comparison, meaning $y_{i^{\prime}}-y_{i}$, but on a difference in probabilities.
@@
 To obtain it, therefore, we need either multiple observations centered on these two profiles or strong assumptions about the behavior of these two profiles.
-In fact, it seems that legal experts require or, at least, expect this kind of comparison (FRA, 2018).
-[^7] We believe that this practice should not come as a surprise to the reader: if it takes more than one throw to check whether a coin is fair, why should it not be the same for a serious accusation like discrimination?
-Intuitively, we wish to control for chance from (or rule out out uncertainty in) the decision process in question to have (some) certainty on the pattern we are testing for.
+Legal experts require or, at least, expect this kind of comparison (FRA, 2018).We believe that this practice should not come as a surprise to the reader.
+For instance, if it takes more than one throw to check whether a coin is fair, why should it not be the same for a serious accusation like discrimination?
+Intuitively, we wish to control for chance from (or rule out out uncertainty in) the decision process in question to have some certainty on the pattern we are testing for.
 This is not possible under a literal comparison.
-Together, these two definitions describe the general setup implicit to all (individual) discrimination tools.
-We provide an overview of these tools in Section 5.
-Given our focus on revisiting the standard comparator, in this work we focus on Definition 2.1.
-Clearly, whatever we derive as the comparator in (2) conditions what outcomes are compared in (3) and, in turn, what we test for and uncover as discrimination when using any discrimination testing tool.
+Together, these two definitions describe the general setup implicit to all discrimination testing tools.
+Given our focus on revisiting the standard CP comparator, in this work we focus on Definition 2.1.
+Clearly, whatever we derive as the comparator following Equation (2) conditions what outcomes are compared in Equation (3) and, in turn, what we test for and uncover as discrimination when using a discrimination testing tool.
 2.2. On (Establishing) Discrimination
-The comparator is intrinsic to how we conceive discrimination.
-This conception is comparative (Westen, 1982; Weerts et al., 2023; Kohler-Hausmann, 2019).
+The comparator is intrinsic to how we conceive discrimination, and this conception is comparative (Westen, 1982; Weerts et al., 2023; Kohler-Hausmann, 2019).
+Now, overall, discrimination in practice can be both a familiar and non-trivial concept.
 We draw a distinction between describing versus proving something as discriminatory, which helps understand discrimination as a social phenomenon (the former) and as a modeling problem (the latter).
@@
 Quoting from Weerts et al.
-(2023), who base their definition on Lippert-Rasmussen (2006): “discrimination can generally be characterized by the morally objectionable practice of subjecting a person (or group of persons) to a treatment in some social dimension that, for no good reason, is disadvantageous compared to the treatment awarded to other persons who are in a similar situation, but who belong to another socially salient group.” A group is considered to be socially salient “if perceived membership of it is important to the structure of social interactions across a wide range of social contexts” (Weerts et al., 2023).
-Defining the socially salient group, e.g., can only be done by agreeing on a specific, shared social context.
-Arriving to this point, e.g., by recognizing certain groups under non-discrimination law, captures the social phenomenon.
-At the core of this view on discrimination there is a clear comparative element.
+(2023), who base their own definition on Lippert-Rasmussen (2006):
+Discrimination can generally be characterized by the morally objectionable practice of subjecting a person (or group of persons) to a treatment in some social dimension that, for no good reason, is disadvantageous compared to the treatment awarded to other persons who are in a similar situation, but who belong to another socially salient group.
+A group is considered to be socially salient “if perceived membership of it is important to the structure of social interactions across a wide range of social contexts” (Weerts et al., 2023).
+Defining the socially salient group can only be done by agreeing on a specific, shared social context.
+Arriving to this point by recognizing certain groups over others under non-discrimination law captures the social phenomenon.
+There is a clear comparative element to this view on discrimination.
 We are not interested in the difference in treatment between any two individuals from different socially salient groups, but in the difference in treatment between two similar (or similarly situated) individuals from different socially salient groups.
-Hence, discrimination occurs when similar (or similarly situated) individuals that differ on membership to a socially salient group are treated differently.
+Discrimination occurs when similar (or similarly situated) individuals that differ on membership to a socially salient group are treated differently.
 Under this setting, discrimination becomes the opposite of equality (Weerts et al., 2023).
@@
 We consider the European Union (EU) context because, one, most fair ML works consider only the US context and, second, the EU—with the AI Act and GDPR—currently leads on the regulation front for algorithms (Álvarez et al., 2024).
-In doing so, we want to show the complex pipeline behind establishing discrimination, and emphasize that it does not begin nor end with Section 2.1.
-Beyond that, although the discussion below concerns EU non-discrimination law, the modeling problem formalized in Section 2.1 is shared among most (Western) conceptions of justice (Westen, 1982) and it is, thus, a robust problem to the legal context.
+In doing so, we want to show the complex pipeline behind establishing discrimination.
+Although the discussion below concerns EU non-discrimination law, the modeling problem formalized in Section 2.1 is shared among most Western conceptions of justice (Westen, 1982) and it is, thus, a robust problem to the legal context.
 There are two forms of discrimination conceived by EU non-discrimination law: direct and indirect discrimination.
@@
 Recent works argue that if the model does not use the protected attribute, but is able to infer it (read, proxy discrimination (Tschantz, 2022)), then it should still fall under direct discrimination (Adams-Prassl et al., 2022).
-  • …“a protected characteristic.” The claim must be based on an attribute protected by non-discrimination law.
+  • …“a protected characteristic.” The claim must be based on an attribute protected by non-discrimination law, such as gender or race.
+These are attributes that summarize historical processes as they were once used, or are still in use, to not only classify individuals but to purposely shape their social and economic opportunities (Sen and Wasow, 2016; Bonilla-Silva, 1997; Michaels and Reed, 2020).
+Non-discrimination law acknowledges the historical injustices toward members of these groups by encoding them as “protected”.
+That is why a discrimination claim is specific to non-discrimination law, meaning it cannot be apply to any characteristic of an individual, only to a protected one.
 This element, on top of the current discussions on intersectional versus multiple discrimination (Xenidis, 2020), is ongoing as there is a fear that models are creating new protected attributes.
@@
 The third element is the one that concerns the comparator as it requires the use of the discrimination testing tools for obtaining the prima facie evidence.
-Proving discrimination, similar to describing discrimination, has a clear comparative element; here, though, it has a modeling angle because the goal of obtaining evidence is explicit.
+Proving discrimination, similar to describing discrimination, has a clear comparative element; here, though, it also has a modeling angle because the goal of obtaining evidence is explicit.
 The idea is still to compare similar (or similarly situated) individuals, though explicitly by defining $d$ in Definition 2.1, materializing what similarity between individuals is within our context.
@@
 The social context that describes discrimination can be and often is taken for granted when modeling the comparator as it is well understood.
-In Example 1.1, e.g., we did not need to motivate females and males as socially salient groups.
+In Example 1.1, for example, we did not need to motivate females and males as socially salient groups.
 The same occurs with most fair ML papers that simply list protected attributes like race and gender to which the proposed methods apply (Ruggieri et al., 2023).
@@
 Discrimination is often conceived as a causal claim on the (in)direct effect of the protected attribute $A$ on the decision outcome of interest $Y$ (Heckman, 1998).
-In Example 1.1, e.g., we are interested in whether Clara did not receive tenure because of her gender?
 Its causal underpinning, which is motivated by non-discrimination law’s own definition of discrimination (Weerts et al., 2023), has long motivated a range of methods for testing discrimination based on counterfactual reasoning.
@@
 This additional emphasize from our side is important when we consider that $i^{\prime}$ can be either a real or fictitious profile, making it hypothetical in the fullest sense.
-That the comparator $i^{\prime}$ can be viewed as a thought experiment is not exploited–both conceptually and experimentally–under the CP comparator.
+Further, that the comparator $i^{\prime}$ can be viewed as a thought experiment is not exploited—both conceptually and experimentally—under the CP comparator.
 We come back to this point in Section 3.
-We diverge from Kohler-Hausmann (2019) precisely because of this emphasize, viewing the dominant counterfactual model of discrimination as an opportunity suited for fair ML, not as a threat.
+This work diverges from Kohler-Hausmann (2019) precisely because of this emphasize, viewing the dominant counterfactual model of discrimination as an opportunity suited for fair ML.
 Why counterfactuals?
@@
 Broadly, the presence of a sufficiently robust pattern suggests a causal relationship, as causality is seen as an invariant force acting on the world.
-[^10] Formally, such view amounts to an interventionist account of causality, which is the basis for both SCM Pearl (2009) and Potential Outcomes Angrist and Pischke (2008), meaning the two most popular causal modeling frameworks.
+[^11] Formally, such view amounts to an interventionist account of causality, which is the basis for both SCM (Pearl (2009)) and Potential Outcomes (Angrist and Pischke (2008)), meaning the two most popular causal modeling frameworks.
 See Woodward (2005) for details.
@@
 If the goal is to detect such pattern(s) and we agree on the premise that discrimination represents a causal pattern, then counterfactual reasoning is the main approach to do so.
-There is a long tradition in causality, especially within SCM, to formulate what constitutes a cause in terms of counterfactual reasoning (Pearl, 2009; Halpern, 2016; Halpern and Pearl, 2001a, b): if intervening $A$ induces a counterfactual outcome $Y^{CF}$ different from the factual outcome $Y$, then $A$ must be a cause of $Y$ [^11] These definitions are much more formal, such as, in terms of probability distributions, but the phrase captures the main idea behind all of them.
+There is a long tradition in causality, especially within SCM, to formulate what constitutes a cause in terms of counterfactual reasoning (Pearl, 2009; Halpern, 2016; Halpern and Pearl, 2001a, b): if intervening $A$ induces a counterfactual outcome $Y^{CF}$ different from the factual outcome $Y$, then $A$ must be a cause of YY.
+[^12] These definitions are much more formal, such as, in terms of probability distributions, but the phrase captures the main idea behind all of them.
 Under this setting, to test for the existence of such pattern would imply to examine the impact of $A$ on the decision-making process by changing $A$ itself.
@@
 Definition 3.1.
-(The MM Comparator) Let $\tilde{d}$ be a protected-aware distance function that, unlike $d$ in (2), considers both $\mathbf{X}$ and $A$ as inputs, capturing the potential influence of $A$ on $\mathbf{X}$.
-For the complainant $i$ $\langle\mathbf{x}_{i},a_{i},y_{i}\rangle$, we define a mutatis mutandis (MM) comparator $i^{\prime}$ as an individual profile $\langle\mathbf{x}_{i^{\prime}},a_{i^{\prime}},y_{i^{\prime}}\rangle$ such that:
+[…]
 (4) $\tilde{d}((\mathbf{x}_{i},a_{i}),(\mathbf{x}_{i^{\prime}},a_{i^{\prime}}))\leq\epsilon\;\;\;\text{and}\;\;\;a_{i}\neq a_{i^{\prime}}$
-where ϵ+italic-$\epsilon\in\mathbb{R}^{+}$ is a threshold.
-Under $\tilde{d}$, two instances with different non-protected attributes, $\mathbf{x}_{i}\neq\mathbf{x}_{i^{\prime}}$, can be considered similar given information, respectively, about $a_{i}$ and $a_{i^{\prime}}$.
-The fact that Definition 3.1 allows for this “dissimilarity” between comparator and complainant motivates the MM denomination.
-The general idea is that under $\tilde{d}$ we should be able, if needed, to account for the influence of $A$ on $\mathbf{X}$ when deriving the $i^{\prime}$ relative to $i$.
-In other words, we “adjust for what needs to be adjusted” as a consequences on “changing” $A$, allowing us to derive an individual profile less conservative than the one conceived under $d$ in Definition 2.1.
-We motivate this argument in the next sub-sections.
-Given the meaning behind $\tilde{d}$, note that, when $a_{i}=a_{i^{\prime}}$, the MM and CP comparators are the same.
-As we move within the same group in $A$, we assume that the influence of $A=a$ on $\mathbf{X}$ is shared among all $i$ individuals with $A=a_{i}$.
-Implicitly, we assume that defining similarity between individuals of the same group in $A$ requires no additional considerations due to shared experience.
-Additionally, when $\mathbf{X}$ and $A$ are independent, the MM and CP comparators are the same even if $a_{i}\neq a_{a^{\prime}}$ as there are no downstream effects of $A$ to $\mathbf{X}$ to be accounted for.
+where $\epsilon\in\mathbb{R}^{+}$ is a threshold. Recall that $i$ must be real profile; $i^{\prime}$ can instead be real or fictitious.
+[…]
+[…]
 3.1. Against Idealized Comparisons
-Let us focus on the standard, CP comparator.
-We argued already that any notion of similarity between individuals implemented through the distance function $d$ in (2) is a normative statement.
-We also argued that, in turn, the comparator derived under $d$ represents the normative statement.
+Let us focus on the standard CP comparator.
+We argued already that any notion of similarity between individuals implemented through the distance function $d(\cdot,\cdot)$ in Equation (2) is a normative statement.
+We also argued that, in turn, the comparator derived under $d(\cdot,\cdot)$ represents the normative statement.
 This is true for the standard comparator as for any other formulation of the comparator, such as the MM comparator.
@@
 First, in Definition 2.1, similarity between individuals comes down to a comparison on the non-protected attributes $\mathbf{X}$ between $i^{\prime}$ and $i$.
-The distance function $d$ is oblivious to (or unaware of) the protected attribute $A$.
-This holds whether $A$ and $\mathbf{X}$ are associated or not as $d$ does not take it into consideration.
+The distance function $d(\cdot,\cdot)$ is oblivious to (or unaware of) the protected attribute $A$.
+This holds whether $A$ and $\mathbf{X}$ are associated or not as $d(\cdot,\cdot)$ does not take it into consideration.
 Second, it follows that the counterfactual representation of the complainant $i$ in the form of the comparator $i^{\prime}$ only varies, ideally, in terms of $A$.
@@
 Is the idealized comparison problematic, though?
-The answer to this questions comes down to having a stance on similarity between individuals, and how it should be operationalized through the $d$ in Definition 2.1.
+The answer to this questions comes down to having a stance on similarity between individuals, and how it should be operationalized through the $d(\cdot,\cdot)$ in Definition 2.1.
 We view it as open to discussion.
-In this work, we do not have a strong stance against the CP comparator as it can be useful, e.g., in uncovering direct discrimination or when an alternative comparator is not possible.
+In this work, we do not have a strong stance against the CP comparator as it can be useful, for example, in uncovering direct discrimination or when an alternative comparator is not possible.
 We do, however, view the discussion against it as necessary for discrimination testing.
-This discussion, e.g., has motivated the MM comparator.
-Key in criticizing the CP comparator and the idealized comparison it represents is Kohler-Hausmann (2019).
+This discussion motivated the MM comparator.
+Pivotal in criticizing the CP comparator and the idealized comparison it represents is Kohler-Hausmann (2019).
 Under the premise that the comparator $i^{\prime}$ in Definition 2.1 represents a counterfactual representation of the complainant $i$, Kohler-Hausmann (2019) criticizes the counterfactual model of discrimination for ignoring how membership to a protected group conditions the material outcomes of individuals and, thus, limits any meaningful counterfactual analysis between protected and non-protected individuals.
@@
 […] [I]f the signifiers of racial categories fundamentally structure the interpretation and relevance of other characteristics or traits of the unit, then it is a mistake to talk about identical units that differ only by raced statues.
-In short, Kohler-Hausmann (2019) implies that the comparator $i^{\prime}$ derived in Definition 2.1 is wrong and, in turn, the tools for testing discrimination built using such comparator are also wrong.
+Kohler-Hausmann (2019) implies that the comparator $i^{\prime}$ derived in Definition 2.1 is wrong and, in turn, the tools for testing discrimination built using such comparator are also wrong.
 [^13] Kohler-Hausmann (2019) later extends the above argument for gender (Hu and Kohler-Hausmann, 2020), covering the two main protected attributes.
@@
 It is not ideal, but it can also improve if we are willing to revisit the derivation of the comparator.
+In Appendix A we discuss additional critical work toward discrimination testing from a non-legal angle.
+Arguments against discrimination testing are not new.
+Similarly, Wachter et al.
+(2020), by classifying fair ML methods as bias preserving or bias transforming depending on how they define similarity, recently pointed out at the limitations of using the conservative CP comparator.
+These arguments, however, are largely absent from the discussion around discrimination testing tools.
+There, the focus is always on tool development, which is to be expected.
+However, as we already argued, the circularity of similarity requires for a deeper considerations from the tool developers.
+3.2. A New Kind of Comparator
+[…]
+[…]
+The MM comparator acts as a target because we, essentially, need to generate it.
+In Example 1.1, for example, why should we consider Vincent when other profiles could also meet the MM comparator definition?
+In other words, we need to have a modeling procedure that motivates the choice of the MM comparator.
+In practice it means estimating the effect of $A$ on $\mathbf{X}$, and adjusting the latter accordingly.
+We discuss this last point in the next section.
+Under this “modeling” interpretation, the proposed MM comparator might seem abstract.
+In fact, why complicate the derivation of the comparator under the MM definition when we already have the CP definition?
+A reason is that, as argued by Wachter et al.
+(2020), non-discrimination law, at least in the EU , aims at substantive equality.
+In practical terms, it means that when testing for discrimination, we are not interested in whether the treatment was the same for similar individuals, but whether the setting (Wachter et al.
+(2020) refers to it as the “status quou”) is suitable for similar individuals to be treated similarly over time.
+This point is similar to the argument on why equal opportunity is better than demographic parity by Hardt et al.
+(2016): there is a difference between hiring equal number of males and female from ensuring that those numbers remain over time.
+In fact, this distinction summarizes why Wachter et al.
+(2020) classifies fair ML definitions into bias preserving (like demographic parity) and bias transforming (like equal opportunity).
+A second reason to consider the MM comparator is that, implicitly, we are already implementing methods that aim at ideal targets.
+We are not proposing a drastic turn in fair ML research.
+A good example of this point is the fairness-accuracy trade-off discussion (Wick et al., 2019; Sharma et al., 2020; Rodolfa et al., 2021; Maity et al., 2020; Dutta et al., 2020).
+Recent works on this topic agree that the trade-off exists as long as the data is biased.
+Hence, demonstrating its existence is trivial given the current state of our society; what is of interest is studying when or under which conditions the trade-off disappears.
+These are hypothetical conditions that, though, serve as future targets to aim for.
+Not surprisingly, in a setting in which $\mathbf{X}$ is non-informative of $A$ or, similarly, $A$ does not influence $\mathbf{X}$, there is no need to consider a trade-off between accuracy and fairness.
+In a fair world, the most accurate model is also the most fair.
+We believe that the MM comparator is meant to capture such world for the purpose of discrimination testing.
+3.3. Extending the Setup
+Let us consider indirect discrimination cases, meaning $A\rightarrow\mathbf{X}\rightarrow Y$.
+Under the causal modeling problem of deriving a comparator, the key question to be answered is: what happens to $\mathbf{X}$ once we manipulate $A$ to derive the comparator?
+Given how Definitions 2.1 and 3.1 differ, in the case of indirect discrimination the CP comparator ignores the downstream effect of $A$ on $\mathbf{X}$ while the MM comparator accounts for it.
+However, while the implementation of the CP comparator is clear, how do we implement the MM comparator?
+Answering this question helps extend the discrimination testing setup in Section 2.1.
+To start, let $g(\cdot)$ represent a generative model such that:
+(5) $\text{for}\;\;\tilde{\mathbf{X}}=g(\mathbf{X},A)\;\;\;\text{we have}\;\;\;\tilde{\mathbf{X}}\perp\!\!\!\!\perp A$
+We view $\tilde{X}$ as an adjusted representation of $\mathbf{X}$ that is removed from the influence of the protected attribute $A$.
+Conceptually, Equation (5) describes the problem of fair representation learning (Zemel et al., 2013) in which we want to learn or have access to $g(\cdot)$ to achieve $\tilde{\mathbf{X}}$.
+Such a model, for example, can be a Variational Auto-Encoder or a Markov Chain (Kingma and Welling, 2014; Salimans et al., 2015; Bengio et al., 2013).
+For our purposes, we view $\tilde{\mathbf{X}}$ as a counterfactual (fair) representation of $\mathbf{X}$ that answers to the question of what would have happened had the individual belonged to the non-protected group instead of the protected group?
+Hence, $g(\cdot)$ represents the counterfactual generation procedure of abduction, action, and prediction described in Section 1.1.
+Given $g(\cdot)$, we should be able to generate the set of non-protected attributes of the complainant $i$’s comparator $i^{\prime}$ and derive the MM comparator $i^{\prime}$ such that:
+(6) $d(\tilde{\mathbf{x}}_{i},\mathbf{x}_{i^{\prime}})\leq\epsilon$
+where, possibly, $\tilde{\mathbf{x}}_{i}\neq\mathbf{x}_{i}$.
+Given the causal relation $A\rightarrow\mathbf{X}$, the comparator $i^{\prime}$ is similar to the counterfactual profile of $i$ but dissimilar to its factual profile.
+Here, $d(\cdot,\cdot)$ represents again a distance function unaware of the protected attribute $A$ in Equation (2).
+Through Equation (6), we summarize the link between the Definitions 2.1 and 3.1 complainant $i$ and its comparator $i^{\prime}$ as:
+(7) $\tilde{d}\big((\mathbf{x}_{i},a_{i}),(\mathbf{x}_{i^{\prime}},a_{i^{\prime}})\big)=d\big(g(\mathbf{x}_{i},a_{i}),\mathbf{x}_{i^{\prime}}\big)$
+where the left-hand-side of the equation represents the general definition of the MM comparator while the right-hand-side an implementation of it under the generative model $g(\cdot)$.
+[…]
+4. An Illustrative Experiment
+To showcase the CP (Definition 2.1) and MM (Definition 3.1) comparators and their impact on discrimination testing, we consider the Law School Admissions example popularized by Kusner et al.
+(2017) and use on it the discrimination tools of situation testing (ST) (Thanh et al., 2011) and counterfactual situation testing (CST) (Álvarez and Ruggieri, 2023, 2025).
+ST and CST are both data mining tools that look for $k$-nearest neighborhoods for complainant-comparator pairs in a dataset.
+We focus on these tools because CST extends ST by implementing the MM comparator over the CP comparator.
+Both tools follow a similar pipeline, even sharing the same distance function for constructing the neighborhoods around the complainant and comparator, with the key difference that ST flips the protected attribute to create its (CP) comparator while CST uses the generated counterfactual of a complainant to create its (MM) comparator.
+CST does so given the SCM $\mathcal{M}$ with DAG $\mathcal{G}$ shown in Figure 1.
+The code for this section is publicly available in a GitHub repository.
+[^14] https://github.com/cc-jalvarez/revisiting-the-comparator.
+The dataset contains $n=21790$ applicants, containing the applicants’ gender ($G$), race ($R$), undergraduate grade-point average ($UGPA$), and law school admissions test score ($LSAT$).
+In terms of the protected attributes, 43.8% of applicants are females and 16.1% are non-whites.
+Following Álvarez and Ruggieri (2023, 2025), we define a decision-maker $b(\cdot)$ that uses only $UGPA$ and $LSAT$ for admitting $Y=1$ or not $Y=0$ an applicant.
+The decision-maker $b(\cdot)$ tests if an applicant’s weighted sum of $UGPA$ and $LSAT$ is above the cutoff $\psi$, where $\psi$ is the median entry requirement of a top US law school.
+[^15] The cutoff is the weighted sum of 60% in $UGPA$ (3.93 over 4.00), and 40% $LSAT$ (46.1 over 48), giving total of 20.8.
+The maximum possible score given $b(\cdot)$ is 22.
+It is based on Yale University Law School: https://www.ilrg.com/rankings/law/index/1/asc/Accept.
+We are interested in testing, separately, whether female and non-white applicants are potentially discriminated under this decision-making process.
+[^16] Intersectionality, meaning what occurs with non-white-female applicants, is not considered under most non-discrimination laws (Xenidis, 2020; Crenshaw, 1989; Álvarez and Ruggieri, 2025).
+Testing-wise, it means that we only consider the claims for gender and race separately.
+CST proceeds as follows.
+Given a tabular dataset of decisions, with each row denoting an individual, it detects whether each protected individual was or not (potentially) discriminated by the decision-maker.
+Each protected individual represents a complainant $i$, and for each $i$ it generates its comparator $i^{\prime}$ by running a counterfactual query on the protected attribute of interest (recall, Section 1.1) based on a SCM.
+Using Gower’s distance (Gower, 1971), CST mines for $k$-similar protected individuals to $i$ and $k$-similar non-protected individuals to $i^{\prime}$.
+The former represents the control group while the latter represents the test group.
+Recall that $i^{\prime}$ does not appear in the input dataset and answers to the question what would have been of $i$ had it been a non-protected individual during the decision process?
+The control group is centered on $i$ while the test group is centered on $i^{\prime}$.
+CST then compares the difference in rejection proportions—meaning how many individuals out of $k$ in a group received a negative decision—between the control and test groups.
+In its strictest form, CST considers as evidence for discrimination if the difference in proportions is strictly greater than zero, meaning the proportion of negative outcomes is greater in the control group compared to the test group.
+Because it evaluates a difference in proportions, CST is able to equip each individual claim with confidence intervals.
+We do not use this aspect of CST here.
+We include the relevant details of CST in Appendix B.
+ST follows similarly, with the key exception that it does not generate the complainant $i^{\prime}$.
+Instead, it builds both control and test groups around $i$ using Gower’s distance (Gower, 1971).
+It means that $i^{\prime}$, in practice, is the most similar non-protected individual to $i$ according to the distance function.
+Both control and test groups are centered around $i$.
+For our purposes in this section, the rest of the procedure is similar to CST.
+Importantly, ST excludes the complainant-comparator pair when looking at the difference in proportion of negative outcomes.
+CST, instead, has the option to include or exclude the “search centers”.
+{tikzpicture}
+[…]
+[…]
+$\mathcal{M}\,$ $\begin{cases}R&\leftarrow U_{R}\\ G&\leftarrow U_{G}\\ UGPA&\leftarrow b_{U}+\beta_{1}\cdot R+\lambda_{1}\cdot G+U_{1}\\ LSAT&\leftarrow b_{L}+\beta_{2}\cdot R+\lambda_{2}\cdot G+U_{2}\end{cases}$
+[…]
+Figure 1.
+The auxiliary causal knowledge with corresponding SCM $\mathcal{M}$ and DAG $\mathcal{G}$ for Law School Admissions (Level 3) (Kusner et al., 2017).
+Let $b_{U}$ and $b_{L}$ denote the intercepts and $\beta_{1}$, $\beta_{2}$, $\lambda_{1}$, $\lambda_{2}$ the regression weights.
+These parameters are estimated from the data using ordinary linear regression.
+We obtain, respectively, the estimates $\hat{b_{U}}=3.21$, $\hat{\beta_{1}}=-0.22$, $\hat{\lambda_{1}}=0.13$ and $\hat{b_{L}}=37.8$, $\hat{\beta_{2}}=-4.64$, $\hat{\lambda_{2}}=-0.61$.
+:
+To illustrate these two tools, consider a female applicant as the complainant $i$.
+ST finds $k$-females and $k$-males with the closest $UGPA$ and $LSAT$ to $i$ and compares the difference in the proportion of rejections between both groups.
+The complainant is considered discriminated if such a difference is significantly different from zero.
+The ideal comparator $i^{\prime}$ in ST is a male with the same $UGPA$ and $LSAT$ as $i$.
+CST does the same with the exception that it first generates $i$’s counterfactual via Figure 1 and uses it to find $k$-males based on the closest $UGPA$ and $LSAT$.
+The female group is constructed in the same way as in ST as the counterfactual generation does not apply to them.
+This is because the counterfactual query of interest (what would have happened to the female applicant had she been male?) is concerned with the group of male applicants.
+There is no ideal comparator $i^{\prime}$ to be found under CST but, instead, a counterfactual representation of $i$ in the form of $i^{\prime}$ with $\tilde{\mathbf{X}}$ (6).
+Table 1. Number (and % over the total cases) of discrimination cases based on race.
+Method $k=25$ $k=50$ $k=100$ $k=200$ $k=500$
+ST 40 (1.14%) 61 (1.74%) 64 (1.83%) 75 (2.14%) 108 (3.08%)
+CST 296 (8.44%) 337 (9.61%) 400 (11.41%) 476 (13.58%) 605 (17.26%)
+Table 2. Number (and % over the total cases) of discrimination cases based on gender.
+Method $k=25$ $k=50$ $k=100$ $k=200$ $k=500$
+ST 117 (1.23%) 229 (2.4%) 258 (2.71%) 368 (3.86%) 492 (5.16%)
+CST 106 (1.11%) 253 (2.65%) 296 (3.1%) 449 (4.71%) 605 (6.34 %)
+Because of how each method constructs the male neighborhoods, ST and CST implement, respectively, the CP and MM comparators.
+Tables 1 and 2 show the number of prima facie discrimination cases identified under ST and CST for, respectively, the protected attributes race (with the goal of answering what would have been the outcome had the applicant been white?) and gender (and what would have been the outcome had the applicant been male?) for different $k$ neighborhood sizes.
+CST generates a counterfactual distribution for each query.
+The CST results are without the complainant-comparator pair.
+This is because ST does not include them.
+We run CST without these “search centers” to have comparable results between both methods.
+The tables illustrate the impact of choosing between the CP and MM comparators, with CST detecting a higher number of prima facie discrimination cases for the same dataset.
+This difference is driven by ST, under the CP comparator, and CST, under the MM comparator, aiming, respectively, for an idealized and a fairness given the difference comparison.
+Intuitively, under a stricter or more conservative notion of similarity, the ST constructs similar neighborhoods based on the protected attribute.
+For instance, we observe this difference in Figure 2 that shows seven randomly chosen individuals considered to have been discriminated by both ST and CST on the basis of their gender.
+We compare the $LSAT$ and $UGPA$ distributions in the “test groups” (i.e., the neighborhoods constructed around the comparator) of ST and CST to the $LSAT$ and $UGPA$’s distribution in the “control group” (i.e., the neighborhood constructed around the complainant), which is the same for both ST and CST.
+The two groups under ST for $LSAT$ and $UGPA$ are similarly distributed while this is not the case under CST.
+Figure 2 illustrates $\tilde{\mathbf{X}}$ (6) and its role in discrimination testing.
+In Figure 3 we carry out the same analysis and find similar results for race.
+The results are starker for race than for gender.
+This point is observed by the difference between ST and CST for each protected attribute across $k$ size.
+For instance, at $k=50$, CST almost triples ST in number of detected cases for race but shows just a marginal increase for gender.
+What is clear from these results is that the effect of race and gender on $LSAT$ and $UGPA$, within at least the context of law school admissions, is not symmetrical—as also confirmed by the parameter estimates reported in the caption of Figure 1.
+As argued in other works (Fryer Jr and Torelli, 2010; Austen-Smith and Fryer, 2005), each protected attribute conditions differently its members as each protected attribute came to be, at least under a social construct perspective (Mallon, 2007), with different purposes.
+Based on these results and our conceptions of the CP and MM comparators, the similar results between ST and CST for gender would indicate a less pronounced impact from gender on $LSAT$ and $UGPA$.
+In other words, not much changes when generating the counterfactual profiles.
+The opposite occurs with race.
+These results seem in line with work on how race can be more detrimental than gender in US college admissions (Baker et al., 2018; Conger and Long, 2013).
+Further, these results point at CST’s capacity to capture less pronounced forms of discrimination.
+Be that as it may, both tables do show evidence of potential discrimination by the decision-maker $b(\cdot)$, with CST detecting more cases.
+As discussed in Section 2.2, detecting discrimination is one aspect of establishing discrimination.
+Indeed, CST paints a worst setting than ST for the admissions process in question, but, regardless, the results of ST or CST would then have to be argued among stakeholders (such as the university itself).
+It varies on a case-by-case basis and the university can always provide, in principle, an objective justification.
+The complainant used by CST for drawing the so-called test groups is generated under counterfactual generation using the abduction, action, and prediction steps (Pearl et al., 2016).
+These profiles determining which male (or white) outcomes are to be compared to the female (or non-white) outcomes are purely hypothetical.
+Each of these generated profiles represents what would have been of the female (or non-white) individual had it been male (or white) in the admissions process to law school.
+In generating them and, in turn, adjusting for the impact of gender (or race) on $UGPA$ and $LSAT$, CST creates individual targets for each protected individual that represent where they would have been in a fairer (read, free from the impact of the protected attribute) world.
+CST tests for equality under these targets.
+These results are specific to ST and CST and the potential limitations inherent to both methods, though they still highlight what it would mean to test for discrimination under the MM comparator instead of the CP comparator.
+Notably, the notion of generating the comparator is non-trivial.
+A method like ST is more conservative in its treatment of similarity between protected and non-protected individuals but, at the same time, it is also more intuitive.
+The CP comparator is the standard because it offers a simple view on similarity.
+A method like CST, instead, enters into a more subjective realm of what similarity should be.
+Although it relies on a modeling counterfactual approach in which the generated counterfactual distribution is unique to the given SCM and the specific intervention, said distribution that support the instances used as comparators is one of other plausible counterfactual representations of how the world should be (Woodward, 2005).
+Some representations are more plausible than others, but still it is not as though CST builds on a shared groundtruth of what the world should be.
+The MM comparator, in that regard, poses a more challenging and, even, uncomfortable challenge of picturing in practice the sort of equality we want to aim for (Wachter et al., 2020).
+Further, in the case of CST, it is worth noting that the derivation of the SCM in Figure 1 is a difficult task.
+For our purposes, which is to showcase the CP versus MM comparison, this limitation is taken for granted.
+However, in practice, it requires either agreeing on the SCM for a given context, or discovering it based on data and knowledge using causal discovery techniques (Vowels et al., 2023; Nogueira et al., 2022).
+Álvarez and Ruggieri (2025) provide a discussion on this limitation, which, in short, can be summarized as a pragmatic view on causal modeling (Loftus, 2024): the SCM help to explicitly state our assumptions about the world.
+That is what makes them useful in setting in which disagreement among stakeholders is likely, such as discrimination testing.
+Additionally, CST depends on a SCM that describes the dataset and the complexity of the dataset used will, thus, condition the SCM.
+Similarly, CST depends on Gower’s distance function.
+Here, we did not explore these limitations.
+These limitations are independent of the MM comparator as a concept, though should be taken into account in case we wish to implement it using SCM.
+Figure 2.
+We show the distributions in the form of box-plots for the $LSAT$ and $UGPA$ for seven randomly chosen complainants based on the protected attribute gender.
+These complainants are found to be discriminated by both ST and CST for $k=100$.
+The so-called control (ctr) are the neighborhoods of similar female profiles to the complainant, which is the same for both tools.
+Similarly, the so-called test groups (tst-st for ST; tst-cf for CST) are the neighborhoods of similar male profiles to the comparator: i.e., the CP comparator for ST, and the MM comparator for CST.
+In both figures we observe a difference between the two kinds of test groups.
+As we would expect, it means that the test and control groups are more similar under ST than under CST.
+Both figures show the impact of the CP and MM comparators.
+:
+Figure 3.
+The same analysis follows as in Figure 2 but based on the protected attribute race, including similar insights and conclusions on the influence of the CP and MM comparators.
+In both figures, we observe a difference between the two kinds of test groups, highlighting again the impact of the CP and MM comparators.
+:
+5. Overview of Discrimination Testing Tools
+In this section, we survey a representative set of tools for discrimination testing focusing on the comparator.
+Different from Section 1.2, the goal is not to position this work but rather to provide an overview of the methods, which span multiple fields, while using our novel formalization of the ceteris paribus and mutatis mutandis comparators.
+We recommend Romei and Ruggieri (2014) for an extensive and multidisciplinary survey.
+All these tools amount to gathering prima facie evidence against a decision-maker, meaning they derive the comparator $i^{\prime}$ for the complainant $i$ and compare their outcomes.
+We view these tools as instances of the counterfactual model of discrimination.
+5.1. Standard Tools
+Here, by standard we mean non-algorithmic tools.
+These tools mainly include natural experiments (e.g., Goldin and Rouse (2000)), field experiments (e.g., Bertrand and Duflo (2017)), and audits (e.g., Fix and Struyk (1993)).
+These tools share a focus on either gathering or evaluating data that is, by design, able to capture the effect of the protected attribute on the decision outcome.
+Such data that is characterized by a before-and-after or a with-and-without mechanism around the protected attribute that ensures the identifiability of the protected attribute’s effect.
+These tools are somewhat outdated for how decisions are taken in modern decision flows, in which ML may enable and even automate key steps.
+Let us start with audits, which are a more holistic kind of tool, very human-dependent, and less focused on modeling.
+Audits consist in examining the decision process in question through human auditors.
+In practice, it means sending a group of experts to observe during a period of time the decision-maker and study its practices to corroborate the claims of the complainant (Fix and Struyk, 1993).
+In the case of the modern decision flow, such a practice only makes sense if the model is used by a human decision-maker or if the focus is to audit the procedure by which the model was trained (Koshiyama et al., 2024), which is increasingly being viewed as a priority over interpretability by regulators (Panigutti et al., 2023).
+The focus is on understanding the overall process and identifying procedural concerns.
+The comparator is not an explicit objective.
+Natural and field experiments, instead, have more of a focus on discrimination as a modeling problem (i.e., identifying the causal effect of the protected attribute) and, thus, prioritize the designing and gathering of data suitable for this purpose.
+Notably, these methods are a product of their time.
+In particular, these methods are a product of the so-called “empirical revolution” experienced by the social sciences in which a mixture of data availability, faster and cheaper computers, and an intuitive causal framework—the Potential Outcomes Framework—made it possible to test causal claims via experiments (Angrist and Pischke, 2008).
+The overall idea behind natural experiments is simple.
+A policy introduction or external shock occurs, creating a before-and-after scenario that, in principle, allows us to test for the effect of the protected attribute.
+We, thus, find ourselves naturally in an experimental setting.
+Goldin and Rouse (2000)’s Orchestrating Impartiality: The Impact of “Blind” Auditions on Female Musicians is a clear example.
+US orchestras before the 1970s used to audition musicians without screens, meaning musicians would perform in full sight to the evaluating committee.
+The US orchestra system back then was male dominated and openly sexist, which motivated the suspicion that there was discrimination against female musicians.
+Orchestras then started to conduct blind auditions to ensure an impartial process.
+Some orchestras even introduced carpets so that female musicians would not be identified by the sound of their heels when walking up the stage.
+Female representation in US orchestras increased after the introduction of the screen.
+Goldin and Rouse (2000), though, using data from the auditions, set out to study whether this amounted as evidence for gender discrimination within the US orchestra system.
+There was indeed a before-and-after mechanism through the introduction of the screen that coincided with an increase in female musicians in orchestras, but to claim a causal effect it was necessary to control for other factors.
+For instance, the introduction of the screen also coincided with an overall increase of female participation in the US work force.
+Goldin and Rouse (2000) used a regression model to control for this and other factors, finding inconclusive evidence for discrimination against female musicians.
+Goldin and Rouse (2000) is a seminal paper, among other things, because it shows how meticulous the model choice has to be to claim discrimination evidence.
+The paper illustrates the importance of deriving a comparator good-enough to isolate the effect of the protected attribute.
+Field experiments, instead, focus on designing an experiment that creates the data for testing the discrimination claim.
+The most famous example is Bertrand and Mullainathan (2004)’s Are Emily and Greg More Employable Than Lakisha and Jamal?
+A Field Experiment on Labor Market Discrimination.
+Also see Bertrand and Duflo (2017) for a survey on field experiments for discrimination.
+Bertrand and Mullainathan (2004) proposed the following experiment: let us create pairs of fictitious CVs tailored to job positions that are the same except for the candidate’s name, where the name is intended to reflect both gender and race within the US context.
+They would create, for example, the same CV for Mike and Jamal (respectively, a common white and black name according to census data at the time); send both CVs to the same job openings via mail (this was done in the early 2000s); and compare which CV received a callback.
+The CV with the black-sounding name was the complainant and the CV with the white-sounding name its comparator.
+All the CVs were drafted by the researchers.
+In the end, they found evidence for discrimination.
+This sort of field experiment is known as a correspondence study.
+See Rooth (2021) for a survey on correspondence studies.
+Another kind of field experiment is situation testing (Rorive, 2009) in which, rather than sending CVs or other forms of correspondence, we send a group of actors that share key physical traits and backstories as the complainant and its comparator.
+Standard tools are at the center of the criticism against the counterfactual model of discrimination.
+This fact is not surprising, as these tools have been around longer, even inspiring the recent algorithmic tools in the next sections.
+They represent the starting point for more novel discrimination testing methods, such as situation testing (Thanh et al., 2011) and counterfactual situation testing (Álvarez and Ruggieri, 2025) both proposing an algorithmic version of the human-based situation testing (Rorive, 2009).
+Heckman (1998), for example, is critical of the effectiveness of these methods in controlling for similarity and other factors and argues that we have been too confident in claiming to be able detect discrimination.
+More recently, as we discussed in earlier sections, Kohler-Hausmann (2019) attacks these methods (she is particularly critical of Bertrand and Mullainathan (2004)) due to their experimental designs and causal treatment of the protected attribute.
+These standard tools rely on the CP comparator.
+As discussed previously, under direct discrimination, these tools are still useful and still have a lot to offer to algorithmic tools in terms of their focus on inferential statistics.
+If, however, we are interested in proving indirect discrimination, then these tools and their experiments need to make a stronger argument for using the CP comparator.
+In particular, this argument should be centered on the kind of signal, as in relevant information, being used by the decision-maker and whether that signal has any link to the protected attribute in question.
+Consider the case of the orchestra auditions (Goldin and Rouse, 2000).
+There the decision-maker cares only for the playing quality of the musician, which should not be related to gender.
+[^17] Though there is a reasonable case to be made that instruments tend to favor male musicians (Criado-Perez, 2019).
+Hence, the signal can be recovered once the screen is imposed as the sounds can go through the screen but all other factors, like the musician’s appearance, are blocked.
+In such a setting, the CP comparator might already be well suited.
+Consider now the case of the CVs (Bertrand and Mullainathan, 2004).
+There, even if we impose a “screen” by removing the names, there is still the possibility that the decision-maker uses signals for candidate potential that are linked to race or gender, such as extracurricular activities or the name of the high-school.
+These are valid signals, or at least signals that could be well argued by the decision-maker as necessary for measuring candidate potential.
+In such a setting, the MM comparator might be better suited.
+5.2. Discrimination Discovery
+The discrimination discovery tools (Kamiran et al., 2013; Ruggieri et al., 2010; Pedreschi et al., 2008) represent the first attempt at what we now call algorithmic fairness.
+These works took the standard methods of the previous section and combined them with data mining methods, shifting the focus from generating or finding the experimental data to exploring data already produced by a decision process.
+This shift in focus was in part field-driven: within data mining and overall knowledge discovery, the focus of the problem formulation is on developing new and efficient algorithms for extracting information from large datasets.
+It is a different focus from the prevalent inferential statistics approach in which the creation of the data is as important as using the most efficient model.
+But this shift also came with a change in how we test for discrimination, especially with the adoption of modern decision flows, as the attention is now on a consistent, algorithmic decision-maker.
+Hence, to test whether the decision-maker discriminated, we could now look at the data it produced.
+The setup in Section 2.1 started to become standard under these tools.
+It was a natural transition from the standard tools as these works borrowed considerably both from non-discrimination law and from the standard methods.
+For example, works by Pedreschi et al.
+(2008) and Ruggieri et al.
+(2010) provided the first formalization to equal opportunity (and, implicitly, to equalized odds) considerably before Hardt et al.
+(2016).
+Kamiran and Calders (2009) did the same for demographic parity.
+Thanh et al.
+(2011) is an illustrative example of the discrimination discovery wave, proposing a $k$-NN implementation of situation testing (Rorive, 2009).
+These tools largely focus on defining the best distance to reach the comparator for a given complainant.
+Some of the tools are closer to the MM comparator, but most of them implement the CP comparator.
+Thanh et al.
+(2011) clearly use the CP comparator as they minimize the distance between the complainant and potential comparators based only on the set of non-protected attributes.
+Zhang et al.
+(2016), instead, offer an extension to (Thanh et al., 2011) via causal weights, focusing on defining a distance function that prioritizes the important non-protected attributes when finding the comparator for a complainant.
+Similarly, Qureshi et al.
+(2020) use a propensity score weight for the distance function with the protected attribute as the treatment.
+None of these tools, however, explicitly address the problem with using the CP comparator or the need to derive the MM comparator.
+For example, there is no discussion on how the protected attribute might influence all other attributes.
+5.3. Algorithmic Fairness
+Discrimination discovery took a broader, more multidisciplinary focus over the past decade and became algorithmic fairness.
+The objectives are the same, but algorithmic fairness is broader in scope, usually aiming to reformulate standard ML problems under a fairness objective (Barocas et al., 2019).
+In turn, many of these tools, even if they are not intended to prove discrimination, are useful for tracking potentially discriminatory decision-making.
+Unsurprisingly, the terms unfair and discriminatory are used interchangeably with algorithmic fairness, sometimes incorrectly (Weerts et al., 2023).
+As discussed in Section 2.2, discrimination claims must align with non-discrimination law in terms of the protected attribute and testing tools represent one aspect of establishing discrimination.
+Given the counterfactual model of discrimination and our focus on the comparator, the main algorithmic fairness works that are central to discrimination testing are counterfactual fairness (Kusner et al., 2017) and individual fairness (Dwork et al., 2012).
+The former establishes the same treatment of the factual and its counterfactual instance under the same decision; the latter establishes the same treatment of similar instances under the same decision.
+Both fairness definitions allude explicitly to the complainant-comparator pair.
+Under the counterfactual model of discrimination, these two concepts intersect as we use the similarity between individuals to approximate the comparison between an individual and its counterfactual-self.
+Works like the FlipTest (Black et al., 2020) or CST (Álvarez and Ruggieri, 2023) build upon these two definitions.
+Overall, this holds for other similar works including (Dutta et al., 2021; Anthis and Veitch, 2023; Salimi et al., 2019).
+Now algorithmic fairness is a vast and ongoing field, with research on fairness definitions and their links to non-discrimination law being one active branch (Wachter et al., 2020, 2021).
+Moving forward, it is important that the CP versus MM comparator discussion is present when looking at (algorithmic) discrimination.
+The choice of method or definition matters beyond providing a complete view when running experiments.
+Such a choice, as discussed, has wider consequences on what we consider discriminatory that are normative concerns rather than modeling concerns.
+Each method, as also argued by Wachter et al.
+(2020), operationalizes a specific ML “intervention” to the systematic unfairness behind the discrimination claim.
+Of this modeling choice we must be aware.
+Moving forward, especially given our formulation of $\tilde{\mathbf{X}}$ in Section 3.3, we view algorithmic fairness tools based on generative models as central to the MM comparator.
+In particular, methods concerned with learning fair representations.
+This research direction started with Zemel et al.
+(2013) and has led to multiple proposed methods over the past decade (e.g., Cerrato et al.
+(2023, 2022); Ganin et al.
+(2016); Louizos et al.
+(2016); Moyer et al.
+(2018); Xie et al.
+(2017); Madras et al.
+(2018)).
+The goal with these methods is to learn a fair representation of the individual, meaning learning a representation of seemingly neutral attributes that removes information from the protected attribute while keeping valuable information of the attributes in question useful for a downstream task.
+We recommend Cerrato et al.
+(2024) for a survey on fair representations.
+Since those complex models tackle the problem (5), they could be, in principle, repurposed for discrimination testing with the goal of deriving an MM comparator.
+Future research is open along this interesting line.
+Our approach relies on SCM, for which such a derivation boils down to counterfactual generation (cfr.
+Section 1.1).
+Algorithmic fairness tools that are more focused on proving discrimination, meaning along the same lines as discrimination discovery, still face the key challenge of linking unfairness to discrimination (Weerts et al., 2023).
+These tools are useful, but it remains to be seen how they fit into the pipeline for establishing discrimination.
+This is not a technical challenge in-itself, but one on whether the technology is embraced or not by the relevant stakeholders.
+Another key challenge, which links directly to the MM comparator, is how these tools address indirect discrimination as most algorithmic decision-makers are not allowed to use the protected attribute as input.
+As illustrated in Section 4, the choice of comparator affects the number of discrimination cases identified.
+These tools, along with the discrimination discovery ones, are capable of generating more complex comparators than the standard tools.
+The question then is whether they should aim or not at implementing the MM comparator over the CP comparator.
+Our view is, at a minimum, they should allow for the possibility of choosing between the two.
+6. Conclusion
+This work discussed the problem of discrimination testing by focusing on the comparator.
+First, we presented the derivation of the comparator as a causal modeling problem.
+It is not the most popular view on discrimination testing (Kohler-Hausmann, 2019), but an important one as it opens impactful opportunities for ML models.
+Second, we introduced the ceteris paribus (CP) comparator and mutatis mutandis (MM) comparator distinction.
+The former refers to the standard comparator and the latter represents a new kind of comparator that departs from idealized comparisons.
+This distinction is novel and will help not only in classifying existing methods but also to better inform new ones and how their choice of similarity between individuals unavoidably impacts normative concerns related to discrimination.
+Additionally, the illustrative experiment based on situation testing (Thanh et al., 2011) and counterfactual situation testing (Álvarez and Ruggieri, 2023, 2025) showcased the impact this distinction has on discrimination testing, motivating the need to consider, at least, the option to choose the kind of comparator when implementing a tool for discrimination testing.
+We focused on causal modeling, particularly, structural causal models (SCM) for generating the MM comparator.
+We stress that said comparator is not tied to this modeling choice, meaning other ML approaches are possible and encouraged.
+Fair representation learning (Cerrato et al., 2024) is the most appealing direction in terms of generative ML for the MM comparator.
+That said, given our causal modeling choice for generating the MM comparator, two limitations are clear and worth discussing.
+First, we use the same procedure as in Kusner et al.
+(2017) to generate the suitable imaginary comparators, meaning using Pearl et al.
+(2016)’s abduction, action, and prediction steps.
+It remains a common procedure that can rely on either a frequentist or Bayesian approach.
+However, it is also a procedure that requires sufficient knowledge about the SCM generating the data, which is complex and potentially non-scalable in practice.
+Other causal ML approaches, such as causal normalizing flows (Javaloy et al., 2023), offer an alternative to generating the counterfactual distribution without having to be explicit about model specifications in the SCM.
+We do not explore these methods as our focus is on the CP versus MM comparison.
+We leave this as future work.
+Second, our extension of the standard discrimination testing setup (Section 3.3) centers on generating an adjusted representation of the non-protected attributes.
+In that sense, it aligns well with the problem of fair representation learning.
+However, an alternative approach could be to define an adjusted similarity function with the same objective.
+We do not consider this approach as, given our use of SCM and overall emphasis on generative ML, the generation of $\tilde{\mathbf{X}}$ fits better.
+We too leave this as future work.
+Exploring different kinds of data-driven distance functions (Blanco-Mallo et al., 2023) as well as searching for similar individuals based on their latent space representation are potential directions to explore.
+Overall, defining a distance function can be as challenging as agreeing on a SCM for generating the MM comparator in the form of $\tilde{\mathbf{X}}$.
+Consider, for instance, the discussions on implementing individual fairness by Dwork et al.
+(2012), which always comes back to the circularity argument in Section 2 inspired by Westen (1982).
+Defining a similarity measure between individuals is inherently contentious.
+At least with the SCM our answer to the what would have been question is answered by the counterfactual distribution, in a way settling the “choice of similarity metric” debate as a generative question.
+Redirecting the focus on a distance function, we believe, would still require to define an imagined objective to learn the distance.
+In fact, said distance could be based on evaluating the factual and counterfactual distributions.
+This is a promising line of research.
+Further, given now the proposed CP and MM comparator distinction, a clear direction for future work is to move beyond an illustrative experiment and survey suitable methods for the MM comparator in a robust experimental setting.
+The goal here was to motivate and formalize such distinction, which is by all means not trivial and often not shared among all of those in the discrimination testing field.
+Future work should move beyond SCM and consider the methods already mentioned as well as other fair ML benchmark datasets to explore the potential limitations of the MM comparator as a modeling problem.
+Finally, this work comes from a ML perspective on discrimination testing.
+We have done our best in incorporating other fields and their views on the need for the MM comparator.
+In doing so, we focused on one aspect—that of providing prima facie evidence—of the full pipeline for establishing discrimination.
+We hope future legal work complements what we have presented here as already the choice of a MM over a CP comparator would expand the current pipeline, requiring additional legal commentary.
+We are certain that the MM comparator and its departure from the CP comparator, both as a conceptual and modeling problem, will motivate further fair ML work on discrimination testing.
+As shown in Example 1.1, it is not straightforward to determine who Clara should be compared to when evaluating her discrimination claim.
+The introduction of the MM comparator helps facilitate this complex and important discussion.
+Acknowledgements.
+We would like to thank Vijay Keswani, Otto Sahlgren, Mattia Cerrato, Alesia Vallenas Coronel, Marcello Di Bello, Nicolò Cangiotti, Francesco Nappo, and Michele Loi for their feedback at different stages of this work.
+This work has received funding from the European Union’s Horizon 2020 research and innovation program for the MSCA project NoBIAS (g.a.
+No.
+860630) and for the project FINDHR (g.a.
+No.
+101070212).
+It reflects only the authors’ views and the European Research Executive Agency is not responsible for any use that may be made of the information it contains.
+Appendix A On The Trouble with Disparity
 Another critical work toward discrimination testing worth mentioning is Michaels and Reed’s article, The Trouble with Disparity (Michaels and Reed, 2020).
@@
 The problem is that, ironically, as Michaels and Reed argue, it has become easier to speak of racial discrimination than income discrimination.
-To be clear, arguments against discrimination testing are not new.
-For instance, Heckman (1998) already criticized ideal comparisons in causal modeling.
-Similarly, Wachter et al.
-(2020), by classifying fair ML methods as bias preserving or bias transforming depending on how they define similarity, recently pointed out at the limitations of using the conservative CP comparator.
-These arguments, however, are largely absent from the discussion around discrimination testing tools.
-There, the focus is always on tool development, which is to be expected.
-However, as we already argued, the circularity of similarity requires for a deeper considerations from the tool developers.
-3.2. A New Kind of Comparator
-Once we allow for a more complex view on the protected attribute $A$ and its influence on the non-protected attributes $\mathbf{X}$, by, e.g., viewing $A$ as a social construct (Kohler-Hausmann, 2019) or acknowledging $A$ as a source of inequality (Michaels and Reed, 2020), the CP comparator appears insufficient.
-Under this view, there is the need to clarify the relationship between $A$ and $\mathbf{X}$, and whether it is in our interest to account for it when testing for discrimination.
-The MM comparator, based on $\tilde{d}$ in (4), is meant to clarify and account for the relationship between $A$ and $\mathbf{X}$, embodying what is refereed to as “fairness given the difference” (Álvarez and Ruggieri, 2023).
-We believe that the MM comparator represents a new kind of comparator, aimed at not only proving discrimination but at shaping the social context behind it.
-The MM comparator represents an ideal target level of equality instead of the formal one captured by the CP comparator due its idealized comparison.
-Conceptually, if $\tilde{d}$ accounts for the influence of $A$ on $\mathbf{X}$, then by comparing $i^{\prime}$ to $i$ when $\mathbf{x}_{i^{\prime}}\neq\mathbf{x}_{i}$ and $a_{i^{\prime}}\neq a_{i}$ we are making a normative statement claiming that these two observably dissimilar individuals should be similar.
-Under counterfactual reasoning, since the counterfactual question is around $A$, the comparator $i^{\prime}$ derived under Definition 3.1 represents where the complainant would have been if free from the effect of being part of the protected group.
-The CP comparator does not allow for this possibility as it only envisions a comparator $i^{\prime}$ all equal but for $A$ to the complainant $i$.
-That is why in Example 1.1, Mike is the CP comparator while Vincent the MM comparator, as Vincent, with his higher number of publications, is meant to represent a Clara that would have carried out her academic career without the social expectations of being the main caregiver of her household.
-The MM comparator acts as a target because we, essentially, need to generate it.
-In Example 1.1, e.g., why should we consider Vincent when other profiles could also meet the MM comparator definition?
-In other words, we need to have a modeling procedure that motivates the choice of the MM comparator.
-In practice it means estimating the effect of $A$ on $\mathbf{X}$, and adjusting the latter accordingly.
-We discuss this last point in the next section.
-Under this “modeling” interpretation, the proposed MM comparator might seem abstract.
-In fact, why complicate the derivation of the comparator under the MM definition when we already have the CP definition?
-A reason is that, as argued by Wachter et al.
-(2020), non-discrimination law, at least in the EU , aims at substantive equality.
-In practical terms, it means that when testing for discrimination, we are not interested in whether the treatment was the same for similar individuals, but whether the setting (Wachter et al.
-(2020) refers to it as the “status quou”) is suitable for similar individuals to be treated similarly over time.
-This point is similar to the argument on why equal opportunity is better than demographic parity by Hardt et al.
-(2016): there is a difference between hiring equal number of males and female from ensuring that those numbers remain over time.
-In fact, this distinction summarizes why Wachter et al.
-(2020) classifies fair ML definitions into bias preserving (like demographic parity) and bias transforming (like equal opportunity).
-A second reason to consider the MM comparator is that, implicitly, we are already implementing methods that aim at ideal targets.
-We are not proposing a drastic turn in fair ML research.
-A good example of this point is the fairness-accuracy trade-off (Wick et al., 2019; Sharma et al., 2020; Rodolfa et al., 2021; Maity et al., 2020; Dutta et al., 2020).
-Recent works on this topic agree that the trade-off exists as long as the data is biased.
-Hence, demonstrating its existence is trivial given the current state of our society; what is of interest is studying when or under which conditions the trade-off disappears.
-These are hypothetical conditions that, though, serve as future targets to aim for.
-Not surprisingly, in a setting in which $\mathbf{X}$ is non-informative of $A$ or, similarly, $A$ does not influence $\mathbf{X}$, there is no need to consider a trade-off between accuracy and fairness.
-In a fair world, the most accurate model is also the most fair.
-We believe that the MM comparator is meant to capture such world for the purpose of discrimination testing.
-3.3. Extending the Setup
-Consider indirect discrimination cases, meaning $A\rightarrow\mathbf{X}\rightarrow Y$ or $Y=f(\mathbf{X})$.
-Under the causal modeling problem of deriving a comparator, the key question to be answered is: what happens to $\mathbf{X}$ once we manipulate $A$ to derive the comparator?
-Given how Definitions 2.1 and 3.1 differ, assuming a case of indirect discrimination, the CP comparator ignores the downstream effect of $A$ on $\mathbf{X}$ while the MM comparator accounts for it.
-However, while the implementation of the former is clear, how do we actually implement the latter?
-How do we define $\tilde{d}$ in (4)?
-Answering this question helps extend the discrimination testing setup in Section 2.1.
-To start, let $g$ represent a generative model such that:
-(5) $\text{for}\;\;\tilde{X}=g(\mathbf{X},A)\;\;\;\text{we have}\;\;\;\tilde{X}\perp\!\!\!\!\perp A$
-We can view, broadly, $\tilde{X}$ as an adjusted representation of $X$ that is removed from the influence of $A$.
-Conceptually, (5) describes the problem of fair representation learning (Zemel et al., 2013) in which we want to learn or have access to $g$ in order to achieve $\tilde{X}$.
-Such a model, e.g., can be a Variational Auto-Encoder or a Markov Chain (Kingma and Welling, 2014; Salimans et al., 2015; Bengio et al., 2013).
-For our purposes, we view $\tilde{X}$ as a counterfactual (fair) representation of $\mathbf{X}$ that answers to the question of what would have happened had the individual belonged to the non-protected group instead of the protected group?
-Hence, $g$ represents the counterfactual generation procedure described in Section 1.1.
-Given $g$, we should be able to generate the set of non-protected attributes of the complainant $i$’s comparator $i^{\prime}$ and derive the MM comparator $i^{\prime}$ such that:
-(6) $d(\tilde{\mathbf{x}}_{i},\mathbf{x}_{i^{\prime}})\leq\epsilon$
-where, possibly, $\tilde{\mathbf{x}}_{i}\neq\mathbf{x}_{i}$.
-Given the causal relation $A\rightarrow\mathbf{X}$, the comparator $i^{\prime}$ is similar to the counterfactual profile of $i$ but dissimilar to its factual profile.
-Here, $d$ represents again a distance function unaware of the protected attribute $A$ in (2).
-Through (6), we summarize the link between the Definitions 2.1 and 3.1 complainant $i$ and its comparator $i^{\prime}$ as:
-(7) $\tilde{d}\big{(}(\mathbf{x}_{i},a_{i}),(\mathbf{x}_{i^{\prime}},a_{i^{\prime}})\big{)}=d\big{(}g(\mathbf{x}_{i},a_{i}),\mathbf{x}_{i^{\prime}}\big{)}$
-where the left-hand-side of the equation represents the general definition of the MM comparator while the right-hand-side an implementation of it under the generative model $g$.
-The relationship (7) allows to derive the MM comparator via $\tilde{d}$ or via $g$ plus $d$.
-Our preference, based on our counterfactual reasoning approach, is with using $g$ plus $d$, which allows for: (i) making it clear the difference between the MM and the CP approaches; (ii) reusing the existing notions of distances and algorithms for inferring them (Kulis, 2013; Ilvento, 2020).
-In that case, the setup in Section 2.1 would require an additional element on top of deriving the comparator and comparing the outcomes: learning $g$, which is of course not limited to counterfactual generation.
-This extended setup for discrimination testing represents an opportunity for ML generative models and a promising path for future work.
-4. An Illustrative Experiment
-[…]
-$\mathcal{M}\,$$\begin{cases}R&\leftarrow U_{R}\\ G&\leftarrow U_{G}\\ UGPA&\leftarrow b_{U}+\beta_{1}\cdot R+\lambda_{1}\cdot G+U_{1},\\ LSAT&\leftarrow\exp\{b_{L}+\beta_{2}\cdot R+\lambda_{2}\cdot G+U_{2}\}\\ \end{cases}$
-$Y$ =$=\begin{cases}1\;\text{if}\;(0.6\cdot UGPA+0.4\cdot LSAT)>\psi\\ 0\;\text{otherwise}\\ \end{cases}$
-Figure 1.
-The causal knowledge with corresponding SCM $\mathcal{M}$ and DAG $\mathcal{G}$ for Law School Admissions.
-Let $b_{U}$ and $b_{L}$ denote the intercepts; $\beta_{1}$, $\beta_{2}$, $\lambda_{1}$, $\lambda_{2}$ the weights; and $UGPA\sim\mathcal{N}$ and $LSAT\sim\text{Poisson}$ the probability distributions.
-To showcase the CP (Definition 2.1) and MM (Definition 3.1) comparators and their impact on discrimination testing, we consider the Law School Admissions example popularized by Kusner et al.
-(2017) and use on it the discrimination tools of situation testing (ST) (Thanh et al., 2011) and counterfactual situation testing (CST) (Álvarez and Ruggieri, 2023).
-ST and CST are data mining tools that look for $k$-nearest neighborhoods for complainant-comparator pairs in a dataset.
-We focus on these two tools because CST extends ST by implementing the MM comparator over the CP comparator.
-Both tools follow a similar pipeline, even sharing the same distance function for constructing the neighborhoods around the complainant and comparator, with the key difference that ST flips the protected attribute to create its (CP) comparator, whilst CST uses the generated counterfactual of a complainant $i$ to create its (MM) comparator $i^{\prime}$.
-The code developed for this section is publicly available in a Github repository.
-[^13] Visit: https://github.com/cc-jalvarez/revisiting-the-comparator.
-The dataset contains $n=21790$ applicants, containing the applicants’ gender ($G$), race ($R$), undergraduate grade-point average ($UGPA$), and law school admissions test score ($LSAT$).
-In terms of the protected attributes, 43.8% of applicants are females and 16.1% are non-whites.
-Following Álvarez and Ruggieri (2023), we define a decision-maker $b$ that uses only $UGPA$ and $LSAT$ for admitting $Y=1$ or not $Y=0$ an applicant.
-The decision-maker $b$ test if a weighted sum of $UGPA$ and $LSAT$ is above a cutoff $\psi$, where $\psi$ is the median entry requirement of a top US law school.
-We are interested in testing, separately, whether female and non-white applicants are potentially discriminated under this decision-making process.
-[^14] Intersectionality, meaning considering what occurs with non-white-female applicants, is not considered under most non-discrimination laws (Xenidis, 2020; Crenshaw, 1989).
-Testing-wise, it means that we can only consider the claims for gender and race separately.
-To illustrate these tools, consider a female applicant as the complainant $i$.
-ST finds $k$-females and $k$-males with the closest $UGPA$ and $LSAT$ to $i$ and compare the difference in the proportion of rejections between both groups.
-The complainant is considered discriminated if such a difference is significantly different from zero at $95\%$ confidence level.
-Under ST, the ideal comparator $i^{\prime}$ in ST is a male with the same $UGPA$ and $LSAT$ as $i$.
-CST does the same with the exception that it first generates $i$’s counterfactual using Figure 1 and uses it to find the $k$ males $UGPA$ and $LSAT$.
-The female group is construct the same way as in ST as the counterfactual generation does not apply to them.
-Under CST, there is no ideal comparator $i^{\prime}$ to be found but, instead, a counterfactual representation of $i$ in the form of $i^{\prime}$ with $\tilde{\mathbf{X}}$ (6).
-Because of how each method constructs the male neighborhoods, ST and CST implement, respectively, the CP and MM comparators.
-Table 1. Number (and % over the total cases) of discrimination cases based on race.
-Method $k=25$ $k=50$ $k=100$ $k=200$ $k=500$
-ST 40 (1.14%) 61 (1.74%) 64 (1.83%) 75 (2.14%) 108 (3.08%)
-CST 296 (8.44%) 337 (9.61%) 400 (11.41%) 476 (13.58%) 605 (17.26%)
-Table 2. Number (and % over the total cases) of discrimination cases based on gender.
-Method $k=25$ $k=50$ $k=100$ $k=200$ $k=500$
-ST 117 (1.23%) 229 (2.4%) 258 (2.71%) 368 (3.86%) 492 (5.16%)
-CST 106 (1.11%) 253 (2.65%) 296 (3.1%) 449 (4.71%) 605 (6.34 %)
-Tables 1 and 2 show the number of discrimination cases identified under ST and CST for, respectively, the protected attributes race (what would have been the outcome had the applicant been white?) and gender (what would have been the outcome had the applicant been male?) for different $k$ neighborhood sizes.
-[^15] The CST results are without the complainant-comparator pair.
-This is because ST does not include them, which allows to compare the results for both methods.
-Results including the pair are similar to those without it.
-See Álvarez and Ruggieri (2023) for details.
-Both tables illustrate the impact of choosing between the CP and MM comparators, with CST detecting a higher number of discrimination cases for the same dataset.
-This difference is driven by ST, under the CP comparator, and CST, under the MM comparator, aiming, respectively for an idealized and a fairness given the difference comparison.
-Intuitively, under a stricter or more conservative notion of similarity, the ST constructs similar neighborhoods based on the protected attribute.
-We observe this in Figure 2 that shows seven randomly chosen individuals considered to have been discriminated by both ST and CST on the basis of their gender.
-We compare the $LSAT$ and $UGPA$ distributions in the “test groups” (i.e., the neighborhoods constructed around the comparator) of ST and CST to the $LSAT$ and $UGPA$’s distribution in the “control group” (i.e., the neighborhood constructed around the complainant), which is the same for both ST and CST.
-The two groups under ST for $LSAT$ and $UGPA$ are similarly distributed while this is not the case under CST.
-Figure 2 illustrates $\tilde{\mathbf{X}}$ (6) and its role in discrimination testing.
-In Figure 3 we carry out the same analysis and find similar results for race.
-Under CST, the complainant used for drawing the so-called test groups is purely generated under counterfactual generation.
-In other words, these profiles determining which male (or white) outcomes are to be compared to the female (or non-white) outcomes are purely hypothetical.
-Each of these generated profiles represents what would have been of the female (or non-white) individual had it been male (or white) in the admissions process to law school.
-In generating them and, in turn, adjusting for the impact of gender (or race) on $UGPA$ and $LSAT$, CST creates individual targets for each protected individual that represent where they would be in a fairer (read, free from the impact of the protected attribute) world.
-CST, thus, tests for equality under these targets.
-Of course, these results are specific to ST and CST and the potential limitations inherent to both methods, though they still highlight what it would mean to test for discrimination under the MM comparator instead of the CP comparator.
-Figure 2.
-We show the distributions in the form of box-plots for the $LSAT$ and $UGPA$ for seven randomly chosen complainants based on the protected attribute gender.
-These complainants are found to be discriminated by both ST and CST for $k=100$.
-The so-called control (ctr) are the neighborhoods of similar female profiles to the complainant, which is the same for both tools.
-Similarly, the so-called test groups (tst-st for ST; tst-cf for CST) are the neighborhoods of similar male profiles to the comparator: i.e., the CP comparator for ST, and the MM comparator for CST.
-In both figures, we observe a difference between the two kinds of test groups.
-As we would expect, it means that the test and control groups are more similar under ST than under CST.
-Figure 3.
-Here, the same analysis follows as in Figure 2 but for the protected attribute race, including similar insights and conclusions on the influence of the CP and MM comparators.
-In both figures, we observe a difference between the two kinds of test groups.
-5. Overview of Discrimination Testing Tools
-In this section, we survey a representative set of tools for discrimination testing, focusing on the comparator.
-We recommend Romei and Ruggieri (2014) for an extensive and multidisciplinary survey.
-All these tools amount to gathering prima facie evidence against a decision-maker, meaning they derive the comparator $i^{\prime}$ for the complainant $i$ and compare their outcomes.
-We view these tools as instances of the counterfactual model of discrimination.
-5.1. Standard Tools
-By standard we mean non-algorithmic tools.
-These tools mainly include natural experiments (e.g., Goldin and Rouse (2000)), field experiments (e.g., Bertrand and Duflo (2017)), and audits (e.g., Fix and Struyk (1993)).
-These tools share a focus on either gathering or evaluating data that is, by design, able to capture the effect of the protected attribute on the decision outcome.
-These tools are somewhat outdated for how decisions are taken in modern decision flow.
-Let us start with audits, which are a more holistic kind of tool, very human-dependent, and less focused on modeling.
-Audits consist in examining the decision process in question through human auditors.
-In practice, it means sending a group of experts to observe during a period of time and study the practices of the decision-maker to corroborate the claims of the complainant (Fix and Struyk, 1993).
-In the case of the modern decision flow, such a practice only makes sense if the model is used by a human decision-maker or if the focus is to audit the procedure by which the model was trained (Koshiyama et al., 2024), which is increasingly being viewed as a priority over interpretability by regulators (Panigutti et al., 2023).
-Natural and field experiments, instead, tend to have more of a focus on discrimination as a modeling problem (i.e., identifying the causal effect of the protected attribute) and, thus, prioritize designing and gathering data suitable for this purpose.
-These methods are a product of their time.
-In particular, these methods are a product of the so-called “empirical revolution” experienced by the social sciences in which a mixture of data availability, better computers, and an intuitive causal framework (the Potential Outcomes Framework) made it possible to test causal claims via experiments (Angrist and Pischke, 2008).
-The overall idea behind natural experiments is simple.
-A policy introduction (or external shock) occurs creating a before-and-after scenario that, in principle, allows us to test for the effect of the protected attribute and find ourselves, naturally, in an experimental setting.
-Goldin and Rouse (2000)’s Orchestrating Impartiality: The Impact of“Blind” Auditions on Female Musicians is a clear example.
-US orchestras before the 1970s used to audition musicians without screens, meaning musicians would perform in full sight to the evaluating committee.
-The US orchestra system back then were male dominated and, some, openly sexist, which motivated the suspicion that there was discrimination against female musicians.
-Orchestras then started to conduct blind auditions to ensure an impartial process.
-Some orchestras even introduced carpets so that female musicians would not be identified by the sound of their heels when walking up the stage.
-Female representation in orchestras increased after the introduction of the screen.
-Goldin and Rouse (2000), though, using data from the auditions, set out to study whether this amounted to evidence for discrimination within the US orchestra system.
-There was indeed a before-and-after mechanism through the introduction of the screen that coincided with an increase in female musicians in orchestras, but to claim a causal effect we needed to control for other factors.
-For instance, the introduction of the screen also coincided with an overall increase of females in the US work force.
-Goldin and Rouse (2000) used a regression model to control for this and other factors, finding inconclusive evidence for discrimination against female musicians.
-Goldin and Rouse (2000) is a seminal paper, among other things, because it shows how meticulous the model choice has to be to claim discrimination evidence.
-It illustrates, implicitly, the importance of deriving a comparator good-enough to isolate the effect of the protected attribute.
-Field experiments, instead, focus on designing an experiment to create the proper data for testing the discrimination claim.
-The most famous example is Bertrand and Mullainathan (2004)’s Are Emily and Greg More Employable Than Lakisha and Jamal?
-A Field Experiment on Labor Market Discrimination.
-Also see Bertrand and Duflo (2017) for a survey on field experiments for discrimination.
-Bertrand and Mullainathan (2004) proposed the following experiment: let us create pairs of fictitious CVs tailored to job positions that are the same except for the candidate’s name, where the name is intended to reflect both gender and race within the US context.
-They would create, e.g., the same CV for Mike and Jamal (respectively, a common white and black name according to census data at the time); send both CVs to the same job openings via mail (this was done in the early 2000s); and compare which CV received a callback.
-The CV with the black-sounding name was the complainant and the CV with the white-sounding name its comparator.
-All the CVs were drafted by the researchers.
-In the end, they found evidence for discrimination.
-This sort of field experiment is known as a correspondence study.
-See Rooth (2021) for a survey on correspondence studies.
-Another kind of field experiment is situation testing (Rorive, 2009) in which, rather than sending CVs or other forms of correspondence, we send a group of actors that share key physical traits and backstories as the complainant and its comparator.
-Standard tools are at the center of the criticism against the counterfactual model of discrimination.
-This fact is not surprising, as these tools have been around longer, even inspiring the recent algorithmic tools.
-Heckman (1998), e.g., is critical of the effectiveness of these methods in controlling for similarity and other factors and argues that we have been too confident in claiming to be able detect discrimination.
-More recently, Kohler-Hausmann (2019) goes after these methods (she is particularly critical of Bertrand and Mullainathan (2004)) due to their experimental designs.
-These standard tools rely on the CP comparator.
-As previously discussed, under direct discrimination, these tools are still useful and have a lot to offer to algorithmic tools in terms of their focus on inferential statistics.
-If, however, we are interested in proving indirect discrimination, then these experiments have to make a strong argument for using the CP comparator.
-In particular, this argument should be centered on the kind of signal (as in relevant information) being used by the decision-maker and whether that signal has any link to the protected attribute.
-Consider the case of the orchestra auditions.
-There, the decision-maker cares only for the playing quality of the musician, which should not be related to gender.
-[^16] Though there is a reasonable case to be made here that instruments tend to favor male musicians (Criado-Perez, 2019).
-Hence, the signal can be recovered once the screen is imposed as the sounds can go through the screen but all other factors, like the musician’s appearance, are blocked.
-Consider now the case of the CVs.
-There, even if we impose a “screen” by removing the names, there is still the possibility that the decision-maker uses signals for candidate potential that are also linked to race or gender, such as extracurricular activities or the name of the high-school.
-These are valid signals, or at least signals that could be well argued by the decision-maker as necessary for measuring candidate potential.
-In such a setting, the MM comparator might be better suited.
-5.2. Discrimination Discovery
-Discrimination discovery tools (Kamiran et al., 2013; Ruggieri et al., 2010; Pedreschi et al., 2008) represent the first attempt at what we now call algorithmic fairness.
-These works took the standard methods and combined them with data mining methods, shifting the focus from generating or finding the experimental data to exploring a data produced by a decision process.
-In part, this shift in focus was field-driven: within data mining and overall knowledge discovery, the focus of the problem formulation is on developing new and efficient algorithms for extracting information from large datasets.
-This is a different focus from the prevalent inferential statics approach in which the creation of the data is as important as using the most efficient model.
-It came with a change in how we test for discrimination, especially with the adoption of modern decision flows, as the attention was now on a consistent, algorithmic decision-maker.
-Hence, to test whether the decision-maker discriminated, we could now look at the data it produced.
-Under these tools, the setup in Section 2.1 started to become standard.
-It was a natural transition.
-Both the decision-maker in question and the tools for evaluating it were scalable.
-These works borrowed considerably both from non-discrimination law and from the standard methods.
-Works by Pedreschi et al.
-(2008) and Ruggieri et al.
-(2010) provided the first formalization to equal opportunity (and, implicitly, to equalized odds) considerably before Hardt et al.
-(2016).
-Kamiran and Calders (2009) did the same for demographic parity.
-Thanh et al.
-(2011) is an illustrative example of the discrimination discovery wave, proposing a $k$-NN implementation of situation testing.
-These tools focus on defining the best distance to reach the comparator for a given complainant.
-Some of them are closer to the MM comparator than to the CP comparator, though.
-Thanh et al.
-(2011) clearly use the CP comparator as they minimize the distance between the complainant and potential comparators based only on the set of non-protected attributes.
-Zhang et al.
-(2016), instead, offer an extension to (Thanh et al., 2011) via causal weights, focusing on defining a distance function that prioritizes the important non-protected attributes when finding the comparator for a complainant.
-Similarly, Qureshi et al.
-(2020) use a propensity score weight for the distance function with the protected attribute as the treatment.
-None of these tools, however, explicitly address the problem with using the CP comparator or the need to derive the MM comparator.
-There is no discussion on how the protected attribute might influence all other attributes.
-5.3. Algorithmic Fairness
-In the past decade, discrimination discovery took a broader, more multidisciplinary focus and became algorithmic fairness.
-The objectives are the same, but algorithmic fairness is broader in scope, usually aiming to reformulate standard ML problems under a fairness objective (Barocas et al., 2019).
-In turn, many of these tools, even if they are not intended to prove discrimination, are useful for tracking potentially discriminatory decision-making.
-Given the counterfactual model of discrimination and our focus on the comparator, the main fairness works useful for discrimination are counterfactual fairness (Kusner et al., 2017) and individual fairness (Dwork et al., 2012).
-The former establishes the same treatment of the factual and its counterfactual instance under the same decision; the latter establishes the same treatment of similar instances under the same decision.
-Under the counterfactual model of discrimination, these two concepts intersect as we use the similarity between individuals to approximate the comparison between an individual and its counterfactual-self.
-Works like the FlipTest (Black et al., 2020) or counterfactual situation testing (Álvarez and Ruggieri, 2023) build upon these two definitions.
-Overall, this holds for other similar works including (Dutta et al., 2021; Anthis and Veitch, 2023; Salimi et al., 2019).
-We view algorithmic fairness tools based on generative models as central to implementing the MM comparator.
-It includes that are not necessarily concerned with discrimination testing but on, e.g., learning fair representations (Zemel et al., 2013).
-Tools that are more focused on proving discrimination, meaning along the same lines as discrimination discovery, still face the key challenge of linking unfairness to discrimination (Weerts et al., 2023).
-These tools are useful, but it remains to be seen how they fit into the pipeline for establishing discrimination.
-This is not a technical challenge in-itself, but one on whether the technology is embraced or not by the relevant stakeholders.
-Another key challenge, which links directly to the MM comparator, is how these tools address indirect discrimination as most algorithmic decision-makers are not allowed to use the protected attribute as input.
-As illustrated in Section 4, the choice of comparator affects the number of discrimination cases identified.
-These tools, along with the discrimination discovery ones, are capable of generating more complex comparators than the standard tools.
-The question then is whether they should aim or not at implementing the MM comparator over the CP comparator.
-Our view is, at a minimum, they should allow for the possibility of choosing between the two.
-6. Conclusion
-We discussed extensively the problem of discrimination testing by focusing on the notion of comparator.
-We presented the derivation of the comparator as a causal modeling problem, emphasizing the role ML methods like generative models can play for better discrimination testing methods.
-We introduced the ceteris paribus (CP) comparator and mutatis mutandis (MM) comparator distinction; the former refers to the standard comparator and the latter represents a new kind of comparator that departs from idealized comparisons.
-Experiments showcase the impact this distinction has on discrimination testing and motivates the need to consider, at least, the option of being able to choose the kind of comparator when implementing a tool for discrimination testing.
-This work comes from a ML perspective on discrimination testing.
-We have done our best in incorporating other fields and their views on the need for the MM comparator.
-Future work should revisit the problem of discrimination testing by considering the full legal procedure for establishing discrimination.
-Future work should consider other generative models for implementing the MM comparator beyond SCM.
-We are certain that the MM comparator and its departure from the CP comparator, both as a conceptual and modeling problem, will motivate further fair ML work on discrimination testing.
-Acknowledgements.
-This work has received funding from the European Union’s Horizon 2020 research and innovation program for the MSCA project NoBIAS (g.a.
-No.
-860630) and for the project FINDHR (g.a.
-No.
-101070212).
-This work reflects only the authors’ views and the European Research Executive Agency is not responsible for any use that may be made of the information it contains.
+Appendix B Supplementary Material for Counterfactual Situation Testing
+In this section, we present additional technical material for CST (Álvarez and Ruggieri, 2023, 2025).
+CST relies on the k-NN algorithm, building the neighborhoods of the control and test groups using a distance function $d(\cdot,\cdot)$ between two individual profiles or tuples.
+Such function only looks at non-protected attributes.
+Let us define the between tuple distance $d(x_{1},x_{2})$ as:
+(8) $d(x_{1},x_{2})=\frac{\sum_{i=1}^{|X|}d_{i}(x_{1,i},x_{2,i})}{|X|}$
+such that $d(x_{1},x_{2})$ averages the sum of the per-attribute distances $d_{i}(x_{1,i},x_{2,i})$ across all attributes in $X$.
+CST handles non-normalized attributes but, as default, we normalize them to insure comparable per-attribute distances.
+The $d_{i}$ used depends on the type of the i-th attribute.
+It equals the overlap measurement ($ol$) if the attribute $X_{i}$ is categorical; otherwise, it equals the normalized Manhattan distance ($md$) if the attribute $X_{i}$ is continuous, ordinal, or interval.
+Hence, $d(\cdot,\cdot)$ amounts to Gower’s distance (Gower, 1971).
+We recall $md$ and $ol$ distances below:
+(9) $md(x_{1,i},x_{2,i})=\frac{|x_{1,i}-x_{2,i}|}{(\max(X_{i})-\min(X_{i}))}$
+(10) $ol(x_{1,i},x_{2,i})=\begin{cases}1&\text{if }x_{1,i}\neq x_{2,i}\\ 0&\text{otherwise}.\end{cases}$
+The main objective of CST is to estimate the difference in proportion of negative decision outcomes for each complainant’s control and test groups:
+(11) $\Delta p=p_{c}-p_{t}$
+with $p_{c}$ and $p_{t}$ represents the count of individuals with a negative decision outcome, respectively, in the control group and test group.
+Recall, that the control group is centered on the complainant $i$ and the test group on its (generated counterfactual) comparator $i^{\prime}$.
+For our purposes, there is potential individual discrimination toward a complainant if $\Delta p>\tau$, where $\tau\in[-1,1]$ represents some accepted deviation.
+Often, $\tau=0$.
+This would amount to a general prima facie discrimination claim.
+CST can also equip such claim with confidence intervals, addressing concerns on whether it is statistically significant or not.
+For more details, we invite the reader to see Álvarez and Ruggieri (2025).